### v1

# S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study

Abstract
In the resource-constrained world of the digital landscape, lightweight cryptography plays a critical role in safeguarding information and ensuring the security of various systems, devices, and communication channels. Its efficient and resource-friendly nature makes it the ideal solution for applications where computational power is limited. In response to the growing need for platform-specific implementations, NIST issued a call for standardization of lightweight cryptography algorithms in 2018. Ascon emerged as the winner of this competition. NIST initally established general evaluation criteria for a standard lightweight scheme including security strength, mitigation against side-channel and fault-injection attacks, and implementation efficiency. To verify the security claims, evaluating the individual components used in any cryptographic algorithm is a crucial step. Substitution box (sbox) is one of the fundamental components used to ensure Shannon’s property of confusion in cryptographic primitives. The quality of a substitution box (S-box) significantly impacts the overall security of a cryptographic primitive. This paper analyzes the S-boxes of six finalists in the NIST Lightweight Cryptography (LWC) standardization process. We evaluate them based on well-established cryptographic properties. Our analysis explores how these properties influence the S-boxes’ resistance against known cryptanalytic attacks and potential implementation-specific vulnerabilities, thus reflecting on their compliance with NIST’s security requirements.
keywords:
Sbox, Lightweight Cryptography, NIST, Substitution Layer
volume: 185issue: 1
S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study
1 Introduction
The growing need for compact, resource-constrained devices performing a variety of functions including but not limited to sensing, identification, and machine control has driven the development of lightweight cryptographic designs and algorithms. The highly evolving yet resource constrained environment raised countless security concerns. As a result, using traditional cryptographic primitives on these platforms while striking an optimum trade-off between compact implementation and security strength became difficult. In order to effectively resolve this issue, NIST lay the foundation for lightweight cryptography standardization in $2017$ [1],[2].
Stemming from the design principles of symmetric cryptography, Lightweight cryptography aims to provide security for Internet of Things (IoT) devices namely sensors, RFID tags, medical implants, smart cards, etc. Design choices such as small block and key size, simple round structure and key schedule and nominal implementation make lightweight ciphers the right choice to ensure a sufficient level of security using less power, computing and memory resources.
As a result of NIST’s call, a total of $57$ proposals were submitted by research teams across the globe [3]. After two rounds of rigorous evaluation process, as described in NIST IR $8369$ [4], $10$ candidates were shortlisted as finalists of the competition, see table 1 In the third and final round, candidates were evaluated on the basis of the claims made in terms of security strength against known attacks ( 112 bits or more); performance on both hardware and software environment; resilience against side channel and fault injection based attacks; and intellectual property rights [5].
As a result, in February $2023$, NIST crowned ASCON family as the new lightweight cryptography standard [6]. The most important evaluation criteria for cryptographic algorithms is measuring their security strength against known attacks such as differential and linear cryptanalysis (and their variants), as well as implementation based attacks. Designers and cryptanalysts across the globe tried their level best to suffice this requirement. As a result, the competition finalists received a noticeable amount of third party security analysis, selective ones highlighted as follows:-
  • Cihangir T. performed key recovery using differential-linear attack on $4$ $5$ rounds of ASCON with a time complexity of $2^{15}$ $2^{33.1}$ [7]. In [8], used undisturbed bits in sbox to perform impossible, truncated and improbable differential attacks on $4$ and $5$ round ASCON. Yanbin L. et al. performed key recovery attack on ASCON $7/12$ round initialization in the nonce misuse scenario with a time complexity of $2^{97}$. They also suggested a forgery attack on $4$, $5$ and $6$ round finalization with timing complexity $2^{9}$, $2^{17}$ $2^{33}$[9]. Maria et al. performed cube based attack on $5$ $6$ rounds with time complexity $2^{35}$ $2^{66}$ and differential-linear key recovery attack on $4$ $5$ rounds of ASCON with a complexity of $2^{18}$ $2^{36}$ [10]. Duarte-Sanchez et al. presented a combined attack using a round reduction hardware trojan to reduce the number of rounds in Ascon to 5 in the initialization phase, and then perform a cube key recovery attack with a timing complexity of $2^{24}$ Further analysis of ASCON includes [11], [12, 13, 14, 15] etc.
  • F.X. et al. show that single trace attacks against leakage resistant PRF used in ISAP can be better mitigated by combining rekeying and shuffling instead of rekeying and masking as suggested in [16]. They also concluded that ISAP is better suited to ensure side channel security on parallelized architectures as compared to low end devices [17].
  • In [18] Yaxin et al. use MILP to find linear characteristics of GIFT-$128$ and then launch a linear cryptanalysis attack on $20$ rounds in order to guess $65$ bits of key. Cao et al. make use of Sun et al. MILP based tool [19] to search for differential characteristics of $7$ $10$ rounds of GIFT-$128$ having a probability of $2^{-15.83}$ $2^{-72.66}$ [20]. Andreas et al. proposed a tool based on bit division property to search for an integral distinguisher on the $11$ rounds of GIFT-$128$ [21]. Wentao et al. used Matsui’s technique to perform related key differential attack on the GIFT-$128$ and locate the best differential trail for upto $14$ rounds [22]. Other attacks include [23, 24, 25, 26] etc.
  • A. Jana and G. Paul proposed two fault models to perform forgery attack on Photon-Beetle in order to first recover the state and then $128$-bit master key. In random fault model, $2^{37.15}$ fault queries, and in known fault model $2^{11.05}$ fault queries are required to recover the entire key [27]. Bart et al. noticed an irregularity in the claimed security bounds leading to a possibility of key recovery attack with $2^{124}$ queries [28]. Other attacks include [29, 30, 31, 32] etc.
  • Haibou Z. et al. proposed interpolation attack on $8$ rounds of Delirium, using $2^{23.1}$ linear equations to perform key recovery with attack complexity $2^{70}$ [33]. Priyanka J. et al. proposed single event fault analysis attack on Dumbo to perform full key recovery using $85$ to $250$ cipher text pairs [34]. A.H. Meraneh et al. theoretically prove the possibility of blind side channel attack on the LFSR used in Elephant. As the internal state of the LFSR is dependant on the secret key. Hence in most use cases this security compromise could lead to secret key being recovered in less than two days [35]. Further attacks include [36, 37].
  • Ahmed A. et al. present an impossible differential attack on $22$ rounds of Skinny-$128$ (underlying cipher in Romulus) [38]. Kazuhiko et al. proposed matching and authenticity attack on Romulus-M with tight security bounds [39]. J. Hua et al. devised a constrained framework to find Meet in the Middle (MITM) characteristics for performing single key recovery attack on $23$ rounds of Skinny-$128$ with time and data complexity of $2^{376}$ $2^{104}$[40]. Other attacks include [41, 42, 43, 44] etc.
Table 1: Methods of ensuring integrity of digital evidence
With S-box Without S-box
ASCON, ISAP, GIFT-COFB,
Photon-Beetle, Elephant, Romulus
Grain-128-AEAD, Sparkle,
Xoodyak, TinyJAMBU
 [^1] Current paper is an extended version of our previous paper [45]
Shannon’s confusion in any symmetric cipher is mainly dependent on the non-linear component used. The most commonly used component to achieve non-linearity in symmetric cryptography is substitution box (sbox). Analysing the cryptographic profile of an sbox plays a vital role in determining the overall strength of a cryptosystem. Literature reports nearly $28$ cryptographic properties [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] that can be used as a tool to estimate the quality of an sbox. Some of the attacks mentioned earlier exploit specific sbox properties in order to devise attack characteristics for a particular cipher. For example, in [34] authors point out that the value of correlation immunity of ASCON sbox makes it susceptible to subset cryptanalysis. However, an entire picture of the cryptographic properties of finalist sboxes is not present in literature and our aim is to fill that gap. Section 2 of this paper elaborates the specifications of six finalists of NIST LWC. Section 3 discusses the fundamental properties of sboxes and classifies them on the basis of relevant attack category. Secondly, upper and lower bounds on these properties are highlighted along with their behaviour under affine equivalence. In Section 4 the results of sbox analysis are discussed in detail. Conclusion is presented in Section 5.
2 Overview of Finalists
This section provides a brief introduction to each of the $6$ finalists from NIST lightweight standardization that make use of sbox as one of the design components.
2.1 ASCON
It is a cryptographic cipher suite presented by Maria et al. [62]. ASCON comprises of three AEAD variants: $128$, $128$ ($a$ and $pq$), and two hashing algorithms: Hash and Hash-$a$. Its main functionality comes from the use of a $320$ bit permutation which is combined with different constants and number of rounds, depending upon the variant. In addition, to achieve non-linearity in the encryption algorithm a $5\times 5$ affine equivalent sbox of $\chi$ mapping from Keccak permutation is used. The ASCON family has been selected by NIST for lightweight standardization because of its security strength and flexibility to be used in multiple applications while bearing a minimum implementation cost [6].
2.2 ISAP
ISAP was proposed by S. Mangard et al. [63], it is a nonce based AEAD cryptographic scheme designed to resist against implementation based attacks at an optimized implementation cost. ISAP uses $400$ bit Keccak and $320$ bit ASCON permutations in its design alongside $5\times 5$ bit ASCON sbox as a non-linear component.
2.3 GIFT-COFB
Proposed by A. Chakraborti. et al.[64]. GIFT-COFB is an encryption scheme that uses combined feedback mode with GIFT-$128$. GIFT-COFB employs $40$ rounds of encryption where the input is: an arbitrary length associated data, an input message of variable length, a nonce of $128$ bits, and a $128$ bit encryption key. Whereas the output is a $128$ bit tag and ciphertext of same length as the input message. A $4\times 4$ bit sbox is used as the confusion component.
2.4 Photon-Beetle
Photon-Beetle is an AEAD based hashing and encryption scheme presented by J. Guo et al.[65]. The encryption routine comprises of $12$ rounds consisting of $4$ separate stages i.e. XOR constant, substitute cells, shift rows and serial mix columns. Photon-Beetle (AEAD) uses the $4$ $4$ bit sbox of Present block cipher in order to achieve non-linearity in its design.
2.5 Elephant
C. Dobarunig et al. [66], proposed Elephant as a permutation based AEAD algorithm. It has $3$ variants based on Spongent and Keccak permutations: Delirium, Jumbo and Dumbo. These encryption variants take $128$ bit key and $96$ bit nonce, however the block sizes are $160$, $176$, $200$. A $4\times 4$ sbox as the non-linear part. In the $80$ rounds of encryption there are $160$ differentially active sboxes.
2.6 Romulus
T. Iwata. et al. [67], proposed Romulus as a nonce based AEAD encryption and hashing scheme that uses Skinny block cipher as underlying encryption algorithm. It operates in $4$ modes of encryption and hashing: nonce misuse based Romulus-M, nonce based Romulus-N, leakage resistant based Romulus-T and hashing based Romulus-H. The encryption algorithm takes $128$ bits of key and nonce, and runs $40$ rounds of encryption. Skinny-$128$ tweakable block cipher [68] uses $8\times 8$ sbox to achieve non-linearity in its design.
Table 2: Classification of s-box properties as per
relevance to existing cryptanalysis techniques
S# Criteria Properties
1 Generic
Balancedness
Permutation
Order of Permutation
Fixed Points (FP)
Opposite Fixed Points (OFP)
Bit Independence Criteria (BIC)
Strict Avalanche Criteria (SAC)
Auto Correlation Table (ACT) [69]
Absolute Indicator (AI) [70]
Sum of Square Indicators (SSI) [70]
2 Linear Cryptanalysis
Nonlinearity (NL) [50, 51]
Linear Approximation Table (LAT) [52, 71]
Linear Approximation Probability (LAP)
Linear Branch Number (LBN)
Linear Structures (LS) [53, 54]
Correlation Immunity (CI) [72]
3 Differential Cryptanalysis
Differential Distribution Table (DDT) [52]
Differential Uniformity (DU) [50, 51]
Differential Branch Number (DBN)
Propagation Characteristics (PC) [73]
Undisturbed Bits [56]
4 Boomerang Cryptanalysis
Boomerang Connectivity Table (BCT)
Boomerang Uniformity (BU) [55]
4 Differential - Linear Cryptanalysis
Linear Structures (LS) [53, 54]
Differential Linear Connectivity Table (DLCT) [57]
Differential Linear Uniformity (DLU) [58]
5 Algebraic Cryptanalysis
Algebraic Degree (AD) [59]
Interpolation Polynomial (IP)
6 Side Channel Analysis
DPA-SNR[60]
Transparency Order (TO) [61]
Table 3: Theoretical Bounds on the Cryptographic Properties of Sboxes
S# Properties
Bounds
For
$4\times 4$ Sbox
For
$5\times 5$ Sbox
For
$8\times 8$ Sbox
Ideal Value
1
$\mathsf{OP}$
$1\leq\mathsf{OP}\leq 2^{n}$ $1\leq\mathsf{OP}\leq 16$ $1\leq\mathsf{OP}\leq 32$ $1\leq\mathsf{OP}\leq 256$ Close to LB
2
$\mathsf{FP}$
$0\leq\mathsf{FP}\leq 2^{n}$ $0\leq\mathsf{FP}\leq 16$ $0\leq\mathsf{FP}\leq 32$ $0\leq\mathsf{FP}\leq 256$ Close to LB
3
$\mathsf{OFP}$
$0\leq\mathsf{OFP}\leq 2^{n}$ $0\leq\mathsf{OFP}\leq 16$ $0\leq\mathsf{OFP}\leq 32$ $0\leq\mathsf{OFP}\leq 256$ Close to LB
4
$\mathsf{BIC}$
$0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ 00
5
$\mathsf{SAC}$
$0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0.5$
6
$\barwedge_{S}$
$\sqrt{{\frac{2^{2n}}{2^{n}-1}}}\leq\barwedge_{S}\leq 2^{n}$
$0<\barwedge_{S}\leq 16$
$0<\barwedge_{S}\leq 32$
$0<\barwedge_{S}\leq 256$
Close to LB
7
$\doublebarwedge_{S}$
$2^{2n}\leq\doublebarwedge_{S}\leq 2^{(3n+m)}$ $256\leq\doublebarwedge_{S}\leq 65536$ $1024\leq\doublebarwedge_{S}\leq 1048576$ $65536\leq\doublebarwedge_{S}\leq 4294967296$ Close to LB
8
$\mathsf{LAT}$
$0\leq\mathsf{LAT}\leq 2^{n-1}$ $0\leq\mathsf{LAT}\leq 8$ $0\leq\mathsf{LAT}\leq 16$ $0\leq\mathsf{LAT}\leq 128$ Close to LB
9
$\xi$
$0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ Close to LB
10
$\mathsf{NL}$
For $n:even$
$0\leq\mathsf{NL}<2^{n-1}-2^{\frac{n}{2}-1}$
For $n:odd$
$0\leq\mathsf{NL}<2^{n-1}-2^{\frac{n-1}{2}}$
$60\leq\mathsf{NL}<$ 6 $120\leq\mathsf{NL}<$ 12 $1200\leq\mathsf{NL}<$ 120 Close to UB
11
$\mathsf{BN_{L}}$
$12\leq\mathsf{BN_{L}}\leq n-$ 1 $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{L}}\leq 4$ $2\leq\mathsf{BN_{L}}\leq 7$ Close to UB
12
$\mathsf{LS}$
$0\leq\mathsf{LS}\leq 2^{n+m}$ $0\leq\mathsf{LS}\leq 256$ $0\leq\mathsf{LS}\leq 1024$ $0\leq\mathsf{LS}\leq 65536$ Close to LB
13
$\mathsf{CI}$
$0\leq\mathsf{CI}\leq n$ $0\leq\mathsf{CI}\leq 4$ $0\leq\mathsf{CI}\leq 5$ $0\leq\mathsf{CI}\leq 8$ Close to LB
14
$\mathsf{DU}$
$2\leq\mathsf{DU}\leq 2^{n}$ $2\leq\mathsf{DU}\leq 16$ $2\leq\mathsf{DU}\leq 32$ $2\leq\mathsf{DU}\leq 256$ Close to LB
15
$\mathsf{BN_{D}}$
$2\leq\mathsf{BN_{D}}\leq\lceil{\frac{2n}{3}}\rceil$ $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{D}}\leq 4$ $2\leq\mathsf{BN_{D}}\leq 6$ Close to UB
16
$\mathsf{PC}$
$0\leq\mathsf{PC}\leq n$  $0\leq\mathsf{PC}\leq 4$ $0\leq\mathsf{PC}\leq 5$ $0\leq\mathsf{PC}\leq 8$ Close to UB
17
$\mathsf{UDB}$
$\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ 00
18
$\beta\mathsf{U}$
$2\leq\beta\mathsf{U}\leq 2^{n}$ $2\leq\beta\mathsf{U}\leq 16$ $2\leq\beta\mathsf{U}\leq 32$ $2\leq\beta\mathsf{U}\leq 256$ Close to LB
19
$\mathsf{DLU}$
$2^{\frac{n}{2}-1}<\mathsf{DLU}\leq 2^{n-1}$ $2<\mathsf{DLU}\leq 8$ $4<\mathsf{DLU}\leq 16$ $8<\mathsf{DLU}\leq 128$ Close to LB
20
$\mathsf{AD}$
$11\leq\mathsf{AD}\leq n-$ 1 $1\leq\mathsf{AD}\leq 3$ $1\leq\mathsf{AD}\leq 4$ $1\leq\mathsf{AD}\leq 7$ Close to UB
21
$\mathsf{DPA_{SNR}}$
For balanced S-box:
$1\leq\mathsf{DPA_{SNR}}\leq 2^{\frac{n}{2}}$
$0\leq\mathsf{DPA_{SNR}}\leq 4$ $60\leq\mathsf{DPA_{SNR}}<$ 6 $0\leq\mathsf{DPA_{SNR}}\leq 16$ Close to LB
22
$\mathsf{TO}$
$0\leq\mathsf{TO}\leq m$ $0\leq\mathsf{TO}\leq 4$ $0\leq\mathsf{TO}\leq 5$ $0\leq\mathsf{TO}\leq 8$ Close to LB
3 Cryptographic Properties of Sboxes and their Classification
The strength of cryptographic primitives is best analysed with the help of well known cryptanalysis techniques. Shannon’s theory of confusion and diffusion plays the foundational role in the design and analysis of iterative ciphers. The aspect of confusion is introduced in symmetric ciphers using highly non-linear components known as sboxes. Designers and cryptanalysts treat these ciphers as a white box, analysing all the underlying mathematical components in order to understand and rule out any potential attack vectors, by imposing design practices based on standard characteristics and theoretically well-defined bounds. Owing to the importance of sboxes in the design of cryptographic algorithms, an extensive research has been done over the years to explain the most refined design and analysis criteria that could be employed in order to secure them against potential attack vectors. These properties can be classified in a number of ways depending whether upon the types of representation in boolean algebra or the statistical and implementation based attacks that could be launched through exploitation of relevant metrics or via studying the inter-dependencies that link one property to another.
This section briefly covers an introduction to around twenty-five cryptographic properties of sboxes that hold significance in the current literature. Each of the properties is related to one or more crytanalysis techniques, that could exploit the non-linear part of a block cipher, see Table 2. In addition, Table 3 presents the upper and lower bounds (UB & LB) on these properties with respect to the dimension of sbox being used [74, 75, 59, 76, 77, 78, 60, 61]. We also talk about the behaviour of these properties under affine equivalence, see Table 4. The properties along with their respective bounds and equivalence characteristics aid designers and cryptanalysts in the selection of good sboxes.
3.1 Diffusion Characteristics
Properties such as balancedness, bijectivity, order of permutation, fixed and opposite fixed points, strict avalanche and bit independence criteria, absolute and sum of square indicator are general pre-requisites that should adhere to defined bounds for any sbox in order to avoid correlations between the input and output bits and serve as a pre-requisite to ensure security against various statistical cryptanalysis attacks.
Balancedness
A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is balanced if it has an equal number of $1$’s and 00’s in its coordinate and component functions. Mathematically;
$P(0)=P(1)=\frac{1}{2}=50\%$ (1)
An imbalanced boolean mapping stays at a risk of achieving high linear approximation probabilities [79]. Balancedness is invariant when under affine equivalence, which means that the number of 00’s ad $1$’s is preserved in $S$ if an affine transformation is applied to it [80, 81, 82].
Permutation
A boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is permutation when the entire input space is utilized in such a way that one input maps only at one output. To fulfill this property it is necessary for $S$ to be a bijective mapping i.e. $n=m$. Permutation property of $S$ is variant under affine transformation [50, 51, 83, 84].
Order of Permutation (OP)
The minimum number of compositions applied to get an identity mapping on the coordinate and component functions of a boolean mapping $S$ is called the order of permutation. OP is not an invariant property under affine equivalence [50, 51, 83, 84].
Fixed Points (FP)
For a vectorial boolean mapping $S$, when an input value $(\varkappa)$ maps upon itself in the output, it is known as fixed point (FP). Mathematically:-
$S(\varkappa)=\varkappa$ (2)
Opposite Fixed Points (OFP)
In a vectorial boolean mapping $S$, if an input value $(\varkappa)$ maps on the inverse of itself in the output, it is known as opposite fixed point (OFP). Mathematically:-
$S(\varkappa)=({\varkappa})^{\prime}$ (3)
If any $S$ has a high number of FPs or OFPs, it is considered to lack the desired randomness so they should kept to a minimum in order to avoid statistical attacks. This can be achieved with the application of an affine transformation, as FP and OFP are affine variant properties [81].
Bit Independence Criteria (BIC)
A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is said to satisfy BIC, if inverting an input bit $u$ has a changing effect on the output bit $v$ i.e. all the avalanche variables $(\gamma_{u},\gamma_{v},....)$ representing difference between bits of input and output of $S$ become pairwise independent [85]. Mathematically, BIC is computed using correlation co-efficient of the avalanche variables:-
$\rho(\gamma_{u},\gamma_{v})=\frac{cov(\gamma_{u},\gamma_{v})}{\sigma(\gamma_{u})\sigma(\gamma_{v})}$ (4)
[…]
Table 4: S-box Properties under Affine Equivalence
Property Affine Variant Affine Invariant
Bijectivity
Balancedness
Permutation
$\mathsf{OP}$
$\mathsf{FP}$
$\mathsf{OFP}$
$\mathsf{BIC}$
$\mathsf{SAC}$
$\mathsf{ACT}$
[…]
$\mathsf{NL}$
$\mathsf{LAT}$
$\mathsf{BN_{L}}$
$\mathsf{LS}$
$\mathsf{CI}$
$\mathsf{DDT}$
$\mathsf{DU}$
$\mathsf{BN_{D}}$
$\mathsf{PC}$
$\mathsf{UDB}$
$\mathsf{BCT}$
$\beta\mathsf{U}$
$\mathsf{DLCT}$
$\mathsf{DLU}$
$\mathsf{AD}$
$\mathsf{IP}$
$\mathsf{DPA-SNR}$
$\mathsf{TO}$
For the coordinate and component functions of $S$ the BIC value should stay in the range of $[0,1]$. An ideal sbox $S$ has a BIC value equal to 00. If an affine transformation is applied to $S$, BIC does not stay invariant [86].
Strict Avalanche Criteria (SAC)
A vectorial boolean mapping $S$ is said to satisfy SAC such that whenever a single input bit $\varkappa$ is inverted, there is a $50$ percent probability of each of the output bit changing. SAC is satisfied if hamming weight $(wt)$ of all the coordinate and component functions of $S$ is $1$. Mathematically, it can be expressed in terms of directional derivative:-
$S_{\gamma}=S(\varkappa)\oplus S(\varkappa\oplus\gamma)$ (5)
Where $\gamma\in F_{2}^{n},\gamma\neq 0,wt(\gamma)=1$ SAC is variant under affine transformation, which means that a function not satisfying SAC could be transformed to a function that satisfies SAC using affine equivalence [87].
Auto Correlation Transform (ACT)
The ACT of $S$ is denoted by $\triangle_{S}(\gamma,\rho)$, where $\gamma\in F_{2}^{n}$, and $\rho\in F_{2}^{m}$. Then ACT of $S$ can be mathematically defined as follows:-
$\triangle_{S}(\gamma,\rho)=\sum_{\varkappa=0}^{2^{n-1}}(-1)^{\rho\cdot(S(\varkappa\oplus\gamma)\oplus S(\varkappa))}$ (6)
ACT of $S$ is invariant under affine equivalence [77].
Absolute Indicator (AI)
The overall impact of auto correlation of a boolean function or a vectorial boolean function is studied by finding the maximum value in ACT spectrum, known as absolute indicator (AI).
Let $f$ be a Boolean function over $F_{2}^{n}$ i.e. $f:F_{2}^{n}\rightarrow F_{2}$. The AI of $f$ is defined as [74]:-
$\barwedge_{f}=\max_{\gamma\in F_{2}^{n}\backslash\{0\}}|\triangle_{f}(\gamma)|$ (7)
Whereas, $\triangle_{f}(\gamma)$ is the autocorrelation of the function $f$ over $F_{2}^{n}$ shifted with $\gamma\in F_{2}^{n}$ and calculated as:-
$F(D_{\gamma}f)=\triangle_{f}(\gamma)=\sum_{\varkappa\in F_{2}^{n}}(-1)^{D_{\gamma}f}$ (8)
Here $D_{\gamma}f=f(x)\oplus f(x\oplus a)$.
For a vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$, using (6) the AI of $S$ is defined as [88]:-
$\barwedge_{S}=\max_{\gamma,\rho\in F_{2}^{n}\backslash\{0\}}|\triangle_{S}(\gamma,\rho)|$ (9)
$\barwedge_{S}$ stays invariant under affine transformation. For balanced vectorial boolean functions, the value of AI is always 00. If a boolean mapping has high $\barwedge_{S}$, it becomes highly susceptible to cube attacks [89, 74].
Sum-of-Squares (SSI)
The SSI of a vectorial boolean mapping $S$ is represented in terms of ACT as follows:-
$\doublebarwedge_{S}=\sum_{\gamma\in F_{2}^{n},,\rho\in F_{2}^{m}}\triangle^{2}_{S}(\gamma,\rho)$ (10)
SSI is not variant under affine equivalence [90]. For achieving good diffusion characteristics in a cryptographic algorithm, sboxes should have low SSI and AI values [91, 74].
3.2 Linear Cryptanalysis
The idea of this technique evolves around studying the non-linear layer of a block cipher to draw a linear approximations profile (LAT) between plaintext, key and the ciphertext to find correlations among them. Matsui introduced non-linearity as a significant criteria in the design of block ciphers [92]. Block ciphers having a highly non-linear layer are resistant towards linear cryptanalysis. To study the number of active sboxes in the linear trails of a block cipher, J. Daemon et al. [93] introduced linear branch number. Theoretically, in order to resist linear (and its variant) cryptanalysis techniques an sbox should have a low linear approximation probability, high non-linearity, high order of linear branch number, equal to none linear structures, and low order correlation immunity.
Linear Approximation Table (LAT)
M. Matsui introduced linear cryptanalysis, as an effective way to define a cryptographic algorithm in terms of linear equations [92, 94, 47]. A linear approximation table (LAT) is constructed by finding the probability of occurrence $P$ of all input ($\varrho$) and output ($\psi$) linear masks where $\varrho,\psi\in F_{2}^{n},F_{2}^{m}$ such that $\varrho\cdot\varkappa=\psi\cdot S(\varkappa)$. LAT can be mathematically represented as follows:-
$\mathsf{LAT}(\varrho,\psi)=\{\#\varkappa\ |\ \varrho\cdot\varkappa=\psi\cdot S(\varkappa)\}$ (11)
Definition 3.1
The maximum LAT table entry is measured in the form of Linear Approximation Probability (LAP), mathematically given as follows:-
$\xi=\frac{1}{2^{n}}\max_{\varrho,\psi\in F_{2}^{n},F_{2}^{m}\backslash\{0\}}\bigg{|}\sum_{\varkappa=1}^{2^{n-1}}\mathsf{LAT}(\varrho,\psi)\bigg{|}$ (12)
The LAT table is of size $2^{n}\times 2^{m}$. The maximum LAT table entry and weight of coefficients in rows and columns stays is invariant under affine transformation, however the distribution of these entries is subject to change. Upper and lower bounds on all the entries of LAT are depicted in Table 3.
Non-linearity (NL)
Non-linearity is the measure of minimum hamming distance between the nonzero component functions of $S$ and their relative affine linear functions. NL of boolean a function $(f)$ is measured as follows:-
$\mathsf{NL_{f}}=\min_{L(\varkappa)\in AL_{n}}dis(f(\varkappa),L(\varkappa))$ (13)
The NL of an $n\times m$ vectorial boolean mapping $S$ is calculated using walsh spectrum $\mathsf{W}(\gamma,\rho)$, where $\gamma,\rho\in F_{2}^{n}$:-
$\mathsf{W}(\gamma,\rho)=\sum_{\varkappa\in F_{2}^{n}}(-1)^{(\gamma\cdot S(\varkappa)\oplus\rho\cdot\varkappa)}$ (14)
$\mathsf{NL_{S}}=2^{n-1}-\frac{1}{2}\max_{\gamma\in F_{2}^{n},\rho\in F_{2}^{m}}|\mathsf{W}(\gamma,\rho)|$ (15)
A high NL value is desirable in order to resist linear cryptanalysis. It is an affine invariant property of any mapping $S$ [95, 96, 97].
Linear Branch Number (LBN)
In block cipher design strategy, branch number is a significant property used to measure the strength of an sbox against linear cryptanalysis. An sbox with high branch number could increase the diffusion power of a cryptographic algorithm. Mathematically:-
$\mathsf{BN_{L}(S)}=\min_{\gamma\in F_{2}^{n},\rho\in F_{2}^{m}\backslash\{0\}}\{w_{t}(\gamma)+w_{t}(\rho)\}$ (16)
LBN is variant under affine transformation i.e. two sboxes can be affine equivalent of each other but they can have different values of LBN [98].
Linear Structures (LS)
J.H. Evertse defines [53]:
A block cipher has Linear Structures (LS) when there exist subsets of input, output and key bits, such that changing all the input and key bits has the same effect on the XOR of the output bits.
Definition 3.2
Let $f:F_{2}^{n}\rightarrow F_{2}$ be a Boolean function, then for some $\gamma\in F_{2}^{n},f(\varkappa\oplus\gamma)\oplus f(\varkappa)=c$ holds for all $\varkappa\in F_{2}^{n}$ where $c\in F_{2}$ is a constant, then $\gamma$ is called a LS of $f(\varkappa)$. More specifically, $\gamma$ is called an invariant LS of $f(\varkappa)$ if $c=0$ and a complementary LS of $f(\varkappa)$ if $c=1$ [91].
Definition 3.3
Let $S:F_{2}^{n}\rightarrow F_{2}^{m}$ be a vectorial boolean mapping if for some $\gamma\in F_{2}^{n},S(\varkappa\oplus\gamma)\oplus S(\varkappa)=c$ holds for all $\varkappa\in F_{2}^{n}$, where $c\in F_{2}^{m}$ is a constant, then $\gamma$ is called a LS of $S(\varkappa)$. More specifically, $\gamma$ is called an invariant LS of $S(\varkappa)$ if $c=0$ and a complementary LS of $S(\varkappa)$ if $c=1$ [99].
Sboxes having LS are vulnerable to attacks more robust than exhaustive key search, such as differential and linear cryptanalysis [100]. LS is an affine variant property of any boolean mapping $S$ [101, 74].
Correlation Immunity (CI)
CI is an independent statistical property of boolean functions depicting linear combinations between the input and output bits of $f$.
Similarly, a boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is said to be $k^{th}$ order correlation immune (given $1<k<n$) , if all its coordinate functions $\gamma\cdot S(\varkappa)$ are independent of the linear combinations in its input. So using (14), we can say that $S$ is $k$-order CI when [102]:-
[…]
CI shows variant behaviour when an affine transformation is applied to $S$ [74]. For a good cryptographic sbox the lower is the order of CI the better. CI shares an inverse relation with NL, hence if CI holds a lower order the value of NL will be high, and vice versa.
3.3 Differential Cryptanalysis
Differential cryptanalysis is a technique that relies on exploiting chosen plaintext pairs and resulting ciphertext pairs, in order to gain the information of secret key. The pairs of plaintext and ciphertext are referred as differentials. And these differentials when combined for multiple rounds become differential characteristics of a cipher under study [47]. The probabilities of occurrence of certain output (ciphertext) differences corresponding fixed input (plaintext) differences are computed in the form of DDT table. A low value of differential uniformity, high differential branch number, high value of propagation characteristics, and no undisturbed bits are a requirement for an sbox to resist against differential cryptanalysis and its variant attacks.
Difference Distribution Table (DDT)
Given two inputs: $\varkappa$ and $\varkappa^{\prime}$ and their corresponding outputs: $\flat$ and $\flat^{\prime}$, the respective input and output differences can be mathematically represented as:-
$\triangle_{\varkappa}=\varkappa\oplus\varkappa^{\prime}$ (18)
$\triangle_{\flat}=\flat\oplus\flat^{\prime}$ (19)
[…]
$\mathsf{DDT(\triangle_{\varkappa},\triangle_{\flat})}=\#\{S(\varkappa\oplus\triangle_{\varkappa})\oplus S(\varkappa)=\triangle_{\flat}\}$ (20)
When an affine transformation is applied to $S$, DDT preserves the maximum entry and weight of coefficients in rows and columns, however their position is subject to a change [103, 104].
Definition 3.4
The maximum achievable value in the DDT of $S$ is called the differential uniformity (DU) of $S$. It can be mathematically represented as follows [105]:-
$\mathsf{DU_{S}}=\max_{\triangle_{\varkappa}\neq 0,\triangle_{\flat}}\#\mathsf{DDT(\triangle_{\varkappa},\triangle_{\flat})}$ (21)
For an sbox to resist differential cryptanalysis a smaller value of DU is required. If an affine transformation is applied to $S$. DU stays invariant [104].
Differential Branch Number (DBN)
The diffusion power of a vectorial boolean mapping $S$ is determined using differential branch number (DBN). It is measured by finding the minimum of hamming distance between two distinct inputs $\varkappa$, $\varkappa^{\prime}$ and their corresponding outputs $S(\varkappa)$, $S(\varkappa^{\prime})$. Mathematically [106]:-
$\mathsf{BN_{D}(S)}=\min_{\varkappa,\varkappa^{\prime}\in F_{2}^{n},\varkappa\neq\varkappa^{\prime}}\{wt(\varkappa\oplus\varkappa^{\prime})\oplus wt(S(\varkappa)\oplus S(\varkappa^{\prime}))\}$ (22)
An sbox having DBN greater than $2$ increases the chances of having more active sboxes in iterative ciphers. DBN does not stay invariant under affine equivalence [98].
Propagation Criteria (PC)
A vectorial boolean function $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is said to satisfy propagation criteria (PC) when $S_{\gamma}=S(\varkappa)\oplus S(\varkappa\oplus\gamma)$ is balanced given $\gamma\in F_{2}^{n}$. Mathematically $S$ satisfy $PC(l)$ iff [74, 91, 107]:-
$\triangle_{S}(\gamma,\rho)=0$ (23)
For $1\leq wt(\gamma)\leq l$.
$PC(l)$ of $S$ is not variant when an affine transformation is applied [87]. An sbox satisfies PC when $ACT_{max}=0$ For an sbox to satisfy SAC, complying with PC is a pre-requisite [74]. Sboxes having low order $PC(l)$ have weak diffusion characteristics [73].
Undisturbed Bits (UDB)
For a specific nonzero input difference $\triangle_{\varkappa}\in F_{2}^{n}$ of an sbox $S:F_{2}^{n}\rightarrow F_{2}^{m}$, some of the bits in the corresponding output difference $\triangle_{\rho}$ remain unchanged, these specific bits are referred to as undisturbed bits (UDB) [108].
Mathematically:-
=$\triangle_{\flat}=\{\triangle_{\rho}=\rho_{j}\in F_{2}^{m}\ |Pr_{S}(\triangle_{\varkappa}\rightarrow\triangle_{\rho})>0\}$ (24)
Here $j=(0,1,2,....m-1)$, if $\rho_{j}=d$, when $d\in F_{2}$ is fixed and $\triangle_{\rho}\in\triangle_{\flat}$ then $S$ has UDBs. UDB is invariant under affine transformation. An sbox should have no UDB in order to avoid the construction of truncated differentials. Currently, there is no theoretically proven upper bound on UDB.
3.4 Boomerang Cryptanalysis
Boomerang attack is a technique of block cipher cryptanalysis based on the similar principle as differential cryptanlysis. However, instead of drawing one differential characteristic for the entire cipher, boomerang attack divides the cipher into two sub ciphers and generates two differentials characteristics also known as boomerang distinguishers. A low value of boomerang uniformity makes an sbox resistant towards boomerang attacks.
Boomerang Connectivity Table (BCT)
The dependency between these two boomerang differentials can have an influence on their combined probability, hence it is of significance to observe the switching effect between the two. In order to achieve this purpose and simplify the analysis, C. Cid et al. proposed the Boomerang Connectivity Table (BCT) [109].
BCT of $S$ is given by a $2^{n}\times 2^{m}$ defined as $\beta T(\triangle_{\varkappa},\nabla_{\flat})$:-
$\begin{split}\beta\mathsf{T}(\triangle_{\varkappa},\nabla_{\flat})=\#\{\varkappa\in F_{2}^{n}\big{|}\ S^{-1}(S(\varkappa)\oplus\nabla_{\flat})\oplus\\ S^{-1}(S(\varkappa\oplus\triangle_{\varkappa})\oplus\nabla_{\flat})=\triangle_{\varkappa}\}\end{split}$ (25)
$\triangle_{\varkappa}\in F_{2}^{n}=$ Input difference of S
$\nabla_{\flat}\in F_{2}^{m}=$ Output difference of S
Boomerang Uniformity (BU)
The highest value in BCT is known as the boomerang uniformity (BU); where $[\triangle_{\varkappa},\nabla_{\flat}]\neq 0$ [55].
$\beta\mathsf{U}=\max_{\triangle_{\varkappa},\nabla_{\flat}\in F_{2}^{n},F_{2}^{m}\backslash\{0\}}\beta T(\triangle_{\varkappa},\nabla_{\flat})$ (26)
BCT and BU of an sbox are invariant under affine transformation [74, 55].
3.5 Differential-Linear Cryptanalysis
Differential linear attack proposed by Hellman and Langford [110] is based on dividing a block cipher into two parts and then combining both differential and linear characteristics into one attack that covers the whole cipher. Hence, the complexity of a differential linear attack is primarily based on the dependency between the two parts.
Differential Linear Connectivity Table (DLCT)
Bar-On et al [57] proposed Differential Linear Connectivity Table (DLCT) in order to study the dependencies between two parts of the cipher so that better differential and linear attack characteristics could be drawn. The DLCT of $S$ is a $2^{n}\times 2^{m}$ table, where rows contain input differences and columns contain output bit masks.
For $\triangle_{\varkappa}\in F_{2}^{n}$ and $\rho\in F_{2}^{m}$, the DLCT entry $(\triangle_{\varkappa},\rho)$ can be computed mathematically as follows:-
$\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho)=\#\{\varkappa\in F_{2}^{n}|\ \rho\cdot S(\varkappa)=\rho\cdot S(\varkappa+\triangle_{\varkappa})\}-2^{n-1}$ (27)
Differential Linear Uniformity (DLU)
The differential linear uniformity (DLU) of $S$ can be defined by taking a maximum on $\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho)$ for all nonzero values of $\triangle_{\varkappa}$ and $\rho$. Mathematically:-
$\Gamma_{S}=\{\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho),\triangle_{\varkappa}\in F_{2}^{n}\backslash\{0\}\mbox{ and }\rho\in F_{2}^{m}\backslash\{0\}\}$ (28)
$\mathsf{DLU}_{S}=\max(\Gamma_{S})$ (29)
Under affine equivalence, DLCT and DLU are invariant [111, 78]. In order to resist against differential linear attack, an sbox should have low DLU.
3.6 Algebraic Cryptanalysis
The basic idea behind algebraic attacks involves converting the entire block cipher round function into a system of equations and then trying to solve these equations in order to deduce the secret key. Both, non-linear and linear components are transformed into polynomials. Unlike the linear component, the number of variables for nonlinear layer increases after every next round. Optimal sboxes are better resistant towards algebraic crypltanalysis if they have a high value of algebraic degree and algebraic immunity, and high degree of interpolation polynomial.
Algebraic Degree (AD)
The polynomial representation of a boolean function $f:F_{2}^{n}\rightarrow F_{2}$, is known as its Algebraic Normal Form (ANF) [74], written as:-
$\begin{split}f(\varkappa)=c_{0}\oplus c_{1}\varkappa_{1}\oplus c_{2}\varkappa_{2}\oplus...\oplus c_{n}\varkappa_{n}\oplus\\ c_{12}\varkappa_{1}\varkappa_{2}\oplus c_{13}\varkappa_{1}\varkappa_{2}\oplus...\oplus c_{(n-1)n}\varkappa_{(n-1)}\varkappa_{n}\\ ...\oplus c_{123.....n}\varkappa_{1}\varkappa_{2}\varkappa_{3}...\varkappa_{n}\end{split}$ (30)
Where the coefficients $c_{i}\in F_{2}$ form the elements of the truth table for ANF of $f(\varkappa)$.
An n-bit vectorial boolean function $S:F_{2}^{n}\rightarrow F_{2}^{m}$ has $f_{j}=f_{1},f_{2},f_{3},.....f_{n}$ coordinate functions and each can be uniquely represented in the form of ANF.
$f_{j}(\varkappa)=\sum_{j=1}^{{n}}C_{j}\prod_{j=1}^{{n}}\varkappa_{j}$ (31)
AD is defined as the number of variables having non-zero coefficients in the highest product term of the ANF. Mathematically, for $S$, AD can be defined as:-
$\mathsf{AD_{S}}=\max_{1\leq j\leq n}{[deg(f_{j}(\varkappa))]}$ (32)
The AD is an affine invariant property of any mapping $S$ [101, 74]. Sboxes having a low AD are vulnerable to higher order differential attacks. AD has an inverse relation with CI, the higher is the order of CI, the lower is AD.
Interpolation Polynomial (IP)
In 1996, Jakobsen and Knudsen introduced interpolation attack on block ciphers[112]. In this attack the round function or sboxes are represented in the form of algebraic expressions. In either case, Lagrange interpolation technique is used to evaluate the algebraic expression. The Lagrange polynomial is constructed using as many plaintext/ ciphertext pairs as the number of unknown coefficients in the algebraic representation. Attack complexity depends on number of terms or degree of the resulting polynomial expression.
For any boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ with input $\varkappa_{i}=\{\varkappa_{1},.....,\varkappa_{n}\}$ and output $\flat_{i}=\{\flat_{1},.....,\flat_{n}\}$ where $i=1,......,n$, we find the Interpolation Polynomial (IP) $P_{i}(\varkappa)$ as follows:
$P_{i}(\varkappa)=\sum_{i=1}^{n}\flat_{i}\prod_{j=1;j\neq i}^{n}\frac{\varkappa-\varkappa_{j}}{\flat_{i}-\flat_{j}}$ (33)
IP is invariant under affine equivalence [113], and a low degree of IP makes any sbox susceptible to interpolation attack.
3.7 Side Channel Analysis
Attacks on cryptographic implementations have evolved significantly since Kocher’s first attempt to use physical leakages from a cryptosystem to extract sensitive information. Due to the increasing threat, designers are always trying to come up with efficient ways to resist against side channel attacks [114]. Starting from the design of components such as sboxes with good cryptographic properties to resist side channels and moving up to the overall cipher level by employing countermeasures such as boolean or polynomial masking. Properties such as DPA SNR and transparency order are nowadays actively used to determine an sbox’s resilience against DPA attacks. However, striking a perfect balance between the metrics related to traditional cryptanalysis and side channel analysis is an open research area. Currently, designers try to achieve cryptographic profiles that are well suited with the requirements of a particular infrastructure or standardization process (as in NIST LWC). In ideal terms, for an sbox to be resistant towards attacks such as DPA, it needs to posses a low DPA SNR and transparency order.
DPA SNR
Differential Power Analysis Signal to Noise Ratio (DPA SNR) was introduced by Sylvain G. et al. in order to improve the quality of leakges from CMOS circuits by modelling ghost peaks [60]. DPA SNR of $S:F_{2}^{n}\rightarrow F_{2}^{m}$ can be mathematically represented in form of walsh spectrum $\mathsf{W}(\gamma,\rho)$ (14), as follows:-
$\mathsf{DPA_{SNR}(S)}=m2^{2n}\left(\sum_{\gamma\in F_{2}^{n}}\left(\sum_{\rho\in F_{2}^{m},wt(\rho)=1}\mathsf{W}(\gamma,\rho)\right)^{4}\right)^{-\frac{1}{2}}$ (34)
As per the authors, DPA SNR increases when resistance of $S$ against linear and differential cryptanalysis increases and vice versa. Thus highlighting that alongside protected implementations, the choice of non-linear component also plays an important role in a block cipher’s resistance against side channel attacks. DPA SNR of $S$ is variant under affine transformation [81].
Transparency Order (TO)
Emmanuel P. et al. extended the concept of (DPA SNR) to evaluate the leakages in a multi-bit DPA setup using hamming weight power model [61].
It is mathematically represented as follows:-
$\begin{split}\mathsf{TO}=\max_{\beta\in F_{2}^{m}}(\big{|}m-2wt(\beta)\big{|})\times\frac{1}{2^{2n}-2^{n}}\times\\ \left(\sum_{\gamma\in F_{2}^{n}}\bigg{|}\sum_{\rho\in F_{2}^{m},wt(u)=1}(-1)^{\rho.\beta}\triangle_{S}(\gamma,\rho)\bigg{|}\right)\end{split}$ (35)
If an sbox has small value of TO, it is considered highly resistant against DPA attacks [61]. When affine transformation is applied to an sbox, TO does not stay invariant [81].
4 Results and Analysis
Table 5 presents the sbox analysis results of Romulus, ASCON/ ISAP, Elephant, Photon-Beetle, and GIFT-COFB. Results of ASCON and ISAP are listed in the same column, since both have same sbox.
Diffusion Metrics
If we look at the basic avalanche characteristics of all these sboxes, they satisfy the criteria for balancedness, ensuring equal number of $1$’s and 00’s in their coordinate and component functions. This helps reduce the chances of getting any high probability linear approximations, in turn assuring the resistance of these sboxes towards linear cryptanalysis[79]. The sboxes under analysis are bijective, so they satisfy the criteria of permutation. From the bounds discussed in Table 3, we have developed an understanding that ideally the sboxes should have a full cycle length for order of permutation. Presumably, sboxes having short iterative periods are vulnerable to attacks so designers take advantage of the affine equivalence criteria and try to improve the order of permutation of sboxes using affine transformation [50, 51, 83, 84]. Out of the six, Ascon/ ISAP sbox has a FP at $\varkappa=255$. Photon-Beetle/ Elephant/ GIFT-COFB, each has one OFPs at $\varkappa=14/8/3$. Sboxes having FPs can alter the overall diffusion power that an iterative cipher has to offer and make it vulnerable to invariant subspace attacks, it can be fixed using affine equivalence. However, in practice, many notable cipher designs avoid the presence of FPs, due to the reason of achieving good results in other affine variant properties. BIC value for all sboxes is $1.0000$ which indicates that the input and output variables have a high correlation amongst them. Furthermore, none of sboxes satisfy SAC. Both these findings indicate that an attacker can gain statistical information about the sboxes and use it to exploit the entire cryptosystem [85, 115]. AI values for all the sboxes meet UB, and the values for that of SSI seem a bit better but do not satisfy the LB requirement. Hence, results from both the properties suggest that sboxes have weak diffusion characteristics.
Linear Metrics
Security towards linear cryptanalysis attacks is ensured when an sbox has a low LAP and high (close to maximum) degree of NL. All $6$ sboxes under analysis have obtained the desired bounds on both properties. LBN of all $6$ sboxes is closer to the LB, which is opposite to the ideal requirements. However, if the design requires it, LBN can be improved using affine transformations. The presence of nonzero LS makes it possible for the cipher to leak significant amount of information if specific group(s) of bits in the input are flipped. All the sboxes are first order correlation immune, which makes them secure against zero-correlation attacks.
Differential and Boomerang Metrics
Almost Perfect Nonlinear (APN) sboxes, in odd dimensions are differentially $2$ uniform [51]. The only odd dimension case in sboxes under study is of Ascon/ ISAP, and since its has a DU: $8$, we can classify it as not an APN. All these sboxes have DU values within the desired bounds. DBN of Elephant and Photon-Beetle satisfies the UB, Ascon/ ISAP is closer to the UB and for Romulus and GIFT-COFB DBN value stands on the LB. A good DBN ensures more number of active sboxes across multiple rounds in block ciphers. As discussed earlier that for an sbox to satisfy SAC one of the pre-requisites is compliance with PC. In Table 5 we can see that all sboxes have PC $=0$, which again points towards the weak avalanche profile.
UDB are detected in all $6$ sboxes. In [108] authors have extensively described how the presence of UDB in an sbox impacts other cryptographic properties, such as the occurrence of LS, as we can see in Table 5. C. Tezcan used these UDB and LS to devise 4 and 5 round impossible, truncated and improbable differentials, and break 5 out of 12 Ascon rounds [116]. In regards to boomerang cryptanalysis, the BU for all sboxes is attaining the UB, which is opposite to the required bounds in order to be secure against Boomerang attacks.
Differential-Linear Metrics
Combined attacks make it easier for an attacker to break a cipher by defining short differential and linear characteristics, as in the case of differential-linear cryptanalysis. Hence, attaining the LB on DLU is crucial to gain resistance against these attacks. In case of sboxes under analysis, we can observe from Table 5, that the DLU is a little way up from the LB, hence paving a way for differential-linear attack with a reduced complexity.
Algebraic Metrics
Previously, we have talked about the inverse relation between CI and AD, and we observed that all $6$ sboxes have a CI $=0$. So given the inverse relation, AD of all sboxes should be obtaining the UB, which in our case is true. AD determines the complexity of an sbox’s ANF, the higher the degree is the more complex it is to launch an algebraic attack on a cipher. Sboxes like those of Ascon/ ISAP having an AD $=2$ are referred to as quadratic and those having AD $=3$ are known as cubic. The functions that posses and AD $=1$ are known as affine.
Side Channel Metrics
During design and analysis phase, DPA SNR and TO of sbox are the two metrics that can be used to determine a cipher’s resistance towards SCA (DPA) attacks. Sboxes that have a high NL and low DU are prone to SCA attacks and have a high resistance towards differential and linear cryptanalysis. Similarly, the sboxes having low values of DPA SNR and TO are resistant towards DPA attacks and vulnerable to differential and linear cryptanalysis. Hence, finding the ideal balance between the two categories of metrics is crucial to the security of a block cipher. Ascon/ ISAP is attaining the almost ideal bounds in terms of NL and DU. Hence, the results of TO and DPA SNR, confirm the notion that Ascon/ ISAP sbox is leaning more on the side of security against linear and differential cryptanalysis. Rest of the sboxes also seem to be attaining UBs on both of SCA metrics.
It is important to note here, that sbox is, although major, yet one mathematical component of a cipher. Designers can always rely on some other design components to balance out the security aspect if using an optimally secure sbox comes at the cost of performance.
Table 5: Results of S-box Analysis
S# Properties Romulus $8\times 8$ Skinny-$128$ Sbox ASCON/ ISAP $5\times 5$ Sbox Elephant $4\times 4$ Sbox GIFT-COFB $4\times 4$ GIFT Sbox Photon-Beetle $4\times 4$ Present Sbox 1 Bijectivity 2 Balancedness 3 Permutation 4 $\mathsf{OP}$ $140$ $26$ $13$ $7$ $9$ 5 $\mathsf{FP}$ $1$ 00 00 00 00 6 $\mathsf{OFP}$ 00 00 $1$ $1$ $1$ 7 $\mathsf{BIC}$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ 8 $\mathsf{SAC}$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ 9 $\barwedge_{S}$ $256$ $32$ $16$ $16$ $256$ 10 $\doublebarwedge_{S}$ $4194304$ $8192$ $1024$ $1024$ $1024$ 11 $\xi$ $0.25000$ $0.25000$ $0.25000$ $0.25000$ $0.25000$ 12 $\mathsf{NL}$ $64$ $8$ $4$ $4$ $4$ 13 $\mathsf{BN_{L}}$ $2$ $3$ $2$ $2$ $2$ 14 $\mathsf{LS}$ $601$ $91$ $9$ $9$ $9$ 15 $\mathsf{CI}$ 00 00 00 00 00 16 $\mathsf{DU}$ $64$ $8$ $4$ $4$ $6$ 17 $\mathsf{BN_{D}}$ $2$ $3$ $3$ $2$ $3$ 18 $\mathsf{PC}$ 00 00 00 00 00 19 $\mathsf{UDB}$ $258$ $35$ $3$ $3$ $6$ 20 $\beta\mathsf{U}$ $16$ $16$ $16$ $16$ $256$ 21 $\mathsf{DLU}$ $\frac{64}{128}=0.5$ $\frac{8}{16}=0.5$ $\frac{4}{8}=0.5$ $\frac{4}{8}=0.5$ $\frac{4}{8}=0.5$ 22 $\mathsf{AD}$ $6$ $2$ $3$ $3$ $3$ 23 $\mathsf{DPASNR}$ $6.312455$ $3.015113$ $2.398501$ $2.128608$ $2.398501$ 24 $\mathsf{TO}$ $7.174510$ $4.258065$ $3.266667$ $3.533333$ $3.466667$
5 Conclusion
In this paper, we have thoroughly discussed $25$ plus well known properties of cryptographic sboxes and described their behaviour under affine transformations. In addition, we have also provided an insight into the minimum and maximum achievable bounds of these properties and the desirable bounds for an optimal sbox.
Our aim was to provide a comprehensive walk-through on the taxonomy of cryptographic properties of sboxes for researchers who are new to the field of design and cryptanalysis of symmetric and lightweight ciphers.
Since the initiation of NIST lightweight cryptography competition, many attacks have been published on different ciphers in the competition. Many of these attacks exploit one of the weakness in the substitution layer. We hope that our analysis of these sboxes and overview on the bounds, will serve as a pre-requisite profile for cryptanalysts to determine new attack distinguishers and propose countermeasures to make the ciphers secure against these attacks.

### v3

@@
-S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study
+
 Abstract
-In the resource-constrained world of the digital landscape, lightweight cryptography plays a critical role in safeguarding information and ensuring the security of various systems, devices, and communication channels.
-Its efficient and resource-friendly nature makes it the ideal solution for applications where computational power is limited.
-In response to the growing need for platform-specific implementations, NIST issued a call for standardization of lightweight cryptography algorithms in 2018.
-Ascon emerged as the winner of this competition.
-NIST initally established general evaluation criteria for a standard lightweight scheme including security strength, mitigation against side-channel and fault-injection attacks, and implementation efficiency.
-To verify the security claims, evaluating the individual components used in any cryptographic algorithm is a crucial step.
-Substitution box (sbox) is one of the fundamental components used to ensure Shannon’s property of confusion in cryptographic primitives.
-The quality of a substitution box (S-box) significantly impacts the overall security of a cryptographic primitive.
-This paper analyzes the S-boxes of six finalists in the NIST Lightweight Cryptography (LWC) standardization process.
-We evaluate them based on well-established cryptographic properties.
-Our analysis explores how these properties influence the S-boxes’ resistance against known cryptanalytic attacks and potential implementation-specific vulnerabilities, thus reflecting on their compliance with NIST’s security requirements.
+Lightweight cryptography was primarily inspired by the design criteria of symmetric cryptography.
+It plays a vital role in ensuring the security, privacy, and reliability of microelectronic devices without compromising the overall functionality and efficiency.
+However, the increasingly platform specific design requirements prompted the development of a standard lightweight algorithm.
+In 2017, NIST put forward security requirements for a standard lightweight scheme — security strength of at least 112 bits against known cryptanalysis attacks, mitigation against side channel and fault injection attacks, and implementation efficiency.
+After three rounds of review, ASCON was crowned as the winner of the competition.
+Evaluating the individual components used in any cryptographic algorithm is an important step in the verification of security claims.
+A fundamental component used to ensure Shannon’s property of confusion in cryptographic primitives is an S-box.
+Hence, the quality of an S-box is a significant contributing factor in the security strength of a cipher.
+In this paper, we evaluate the S-boxes of 6 NIST LWC competition finalists based on well-known cryptographic properties, and comment on how the results reflect upon NIST security requirements.
+Our findings have revealed that these S-boxes do not comply with the basic notions of avalanche, making it vulnerable to high-order sophisticated cryptanalysis.
 keywords:
-Sbox, Lightweight Cryptography, NIST, Substitution Layer
-volume: 185issue: 1
-S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study
+NIST, Lightweight Cryptography, Substitution layer, Cryptanalysis, S-box
 1 Introduction
-The growing need for compact, resource-constrained devices performing a variety of functions including but not limited to sensing, identification, and machine control has driven the development of lightweight cryptographic designs and algorithms.
-The highly evolving yet resource constrained environment raised countless security concerns.
-As a result, using traditional cryptographic primitives on these platforms while striking an optimum trade-off between compact implementation and security strength became difficult.
-In order to effectively resolve this issue, NIST lay the foundation for lightweight cryptography standardization in $2017$ [1],[2].
-Stemming from the design principles of symmetric cryptography, Lightweight cryptography aims to provide security for Internet of Things (IoT) devices namely sensors, RFID tags, medical implants, smart cards, etc.
-Design choices such as small block and key size, simple round structure and key schedule and nominal implementation make lightweight ciphers the right choice to ensure a sufficient level of security using less power, computing and memory resources.
-As a result of NIST’s call, a total of $57$ proposals were submitted by research teams across the globe [3].
-After two rounds of rigorous evaluation process, as described in NIST IR $8369$ [4], $10$ candidates were shortlisted as finalists of the competition, see table 1 In the third and final round, candidates were evaluated on the basis of the claims made in terms of security strength against known attacks ( 112 bits or more); performance on both hardware and software environment; resilience against side channel and fault injection based attacks; and intellectual property rights [5].
-As a result, in February $2023$, NIST crowned ASCON family as the new lightweight cryptography standard [6].
+Application specific computing has become an emerging trend in modern information technology due to the wide usability of smart devices.
+The highly evolving yet resource constrained environment enumerates to countless security concerns.
+It is difficult to use traditional cryptographic primitives on these platforms while striking an optimal trade-off between compact implementation and security strength [1].
+In order to effectively resolve this issue, NIST laid the foundation for lightweight cryptography standardization in $2017$ [2].
+Lightweight cryptography aims to provide security for Internet of Things (IoT) devices, namely sensors, RFID tags, medical implants, smart cards, etc.
+Design choices such as small block and key size, simple round structure and key schedule, and nominal implementation make lightweight ciphers the right choice to ensure a sufficient level of security using less power, computing, and memory resources.
+As a result of NIST’s call, a total of $57$ proposals were submitted by research teams across the globe.
+After two rounds of the rigorous evaluation process, as described in NIST IR $8369$ [3], $10$ candidates were shortlisted as finalists of the competition (see Table 1 [4]).
+In the third and final round, candidates were evaluated on the basis of claims made in terms of security strength against known attacks ($112$ bits or more); performance in both hardware and software environments; resilience against side channel and fault injection attacks; and intellectual property rights [5].
+In February $2023$, NIST declared the ASCON family as the new lightweight cryptography standard [6].
 The most important evaluation criteria for cryptographic algorithms is measuring their security strength against known attacks such as differential and linear cryptanalysis (and their variants), as well as implementation based attacks.
 Designers and cryptanalysts across the globe tried their level best to suffice this requirement.
-As a result, the competition finalists received a noticeable amount of third party security analysis, selective ones highlighted as follows:-
-  • Cihangir T.
-performed key recovery using differential-linear attack on $4$ $5$ rounds of ASCON with a time complexity of $2^{15}$ $2^{33.1}$ [7].
-In [8], used undisturbed bits in sbox to perform impossible, truncated and improbable differential attacks on $4$ and $5$ round ASCON.
-Yanbin L.
-et al.
-performed key recovery attack on ASCON $7/12$ round initialization in the nonce misuse scenario with a time complexity of $2^{97}$.
-They also suggested a forgery attack on $4$, $5$ and $6$ round finalization with timing complexity $2^{9}$, $2^{17}$ $2^{33}$[9].
-Maria et al.
-performed cube based attack on $5$ $6$ rounds with time complexity $2^{35}$ $2^{66}$ and differential-linear key recovery attack on $4$ $5$ rounds of ASCON with a complexity of $2^{18}$ $2^{36}$ [10].
-Duarte-Sanchez et al.
-presented a combined attack using a round reduction hardware trojan to reduce the number of rounds in Ascon to 5 in the initialization phase, and then perform a cube key recovery attack with a timing complexity of $2^{24}$ Further analysis of ASCON includes [11], [12, 13, 14, 15] etc.
-  • F.X.
-et al.
-show that single trace attacks against leakage resistant PRF used in ISAP can be better mitigated by combining rekeying and shuffling instead of rekeying and masking as suggested in [16].
-They also concluded that ISAP is better suited to ensure side channel security on parallelized architectures as compared to low end devices [17].
-  • In [18] Yaxin et al.
-use MILP to find linear characteristics of GIFT-$128$ and then launch a linear cryptanalysis attack on $20$ rounds in order to guess $65$ bits of key.
-Cao et al.
-make use of Sun et al.
-MILP based tool [19] to search for differential characteristics of $7$ $10$ rounds of GIFT-$128$ having a probability of $2^{-15.83}$ $2^{-72.66}$ [20].
-Andreas et al.
-proposed a tool based on bit division property to search for an integral distinguisher on the $11$ rounds of GIFT-$128$ [21].
-Wentao et al.
-used Matsui’s technique to perform related key differential attack on the GIFT-$128$ and locate the best differential trail for upto $14$ rounds [22].
-Other attacks include [23, 24, 25, 26] etc.
-  • A.
-Jana and G.
-Paul proposed two fault models to perform forgery attack on Photon-Beetle in order to first recover the state and then $128$-bit master key.
-In random fault model, $2^{37.15}$ fault queries, and in known fault model $2^{11.05}$ fault queries are required to recover the entire key [27].
-Bart et al.
-noticed an irregularity in the claimed security bounds leading to a possibility of key recovery attack with $2^{124}$ queries [28].
-Other attacks include [29, 30, 31, 32] etc.
-  • Haibou Z.
-et al.
-proposed interpolation attack on $8$ rounds of Delirium, using $2^{23.1}$ linear equations to perform key recovery with attack complexity $2^{70}$ [33].
-Priyanka J.
-et al.
-proposed single event fault analysis attack on Dumbo to perform full key recovery using $85$ to $250$ cipher text pairs [34].
-A.H.
-Meraneh et al.
-theoretically prove the possibility of blind side channel attack on the LFSR used in Elephant.
-As the internal state of the LFSR is dependant on the secret key.
-Hence in most use cases this security compromise could lead to secret key being recovered in less than two days [35].
-Further attacks include [36, 37].
-  • Ahmed A.
-et al.
-present an impossible differential attack on $22$ rounds of Skinny-$128$ (underlying cipher in Romulus) [38].
-Kazuhiko et al.
-proposed matching and authenticity attack on Romulus-M with tight security bounds [39].
-J.
-Hua et al.
-devised a constrained framework to find Meet in the Middle (MITM) characteristics for performing single key recovery attack on $23$ rounds of Skinny-$128$ with time and data complexity of $2^{376}$ $2^{104}$[40].
-Other attacks include [41, 42, 43, 44] etc.
-Table 1: Methods of ensuring integrity of digital evidence
+As a result, the competition finalists received a noticeable amount of third-party security analysis [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], a few highlighted in Table 2:-
+Table 1: Classification of NIST LWC Finalists based on Presence/Absence of S-box
 With S-box Without S-box
@@
 Xoodyak, TinyJAMBU
- [^1] Current paper is an extended version of our previous paper [45]
-Shannon’s confusion in any symmetric cipher is mainly dependent on the non-linear component used.
-The most commonly used component to achieve non-linearity in symmetric cryptography is substitution box (sbox).
-Analysing the cryptographic profile of an sbox plays a vital role in determining the overall strength of a cryptosystem.
-Literature reports nearly $28$ cryptographic properties [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] that can be used as a tool to estimate the quality of an sbox.
-Some of the attacks mentioned earlier exploit specific sbox properties in order to devise attack characteristics for a particular cipher.
-For example, in [34] authors point out that the value of correlation immunity of ASCON sbox makes it susceptible to subset cryptanalysis.
-However, an entire picture of the cryptographic properties of finalist sboxes is not present in literature and our aim is to fill that gap.
+Table 2: Summary of Third-Party Security Analysis on NIST LWC Finalists
+Cipher Attack Type Rounds Complexity Reference
+ASCON Differential-Linear 44 55 Time $2^{15}$, $2^{18}$ Time $2^{33.1}$, $2^{36}$ [7] [9]
+Impossible Differential Truncated Differential $4,5$ $4,5$ Data $2^{2}$, $2^{109}$ Data $2^{2}$, $2^{109}$ [45]
+Cube Attack 55 66 Time $2^{35}$ Time $2^{66}$ [9]
+Key Recovery (Nonce Misuse) Init ($7/12$) $8,9$ Time $2^{97}$, $2^{101}$, $2^{123.92}$ [8] [46]
+Forgery Attack Final (46)(4-6) Time $2^{9}$, $2^{17}$ & $2^{33}$ [8]
+Round Reduction ++ Cube Init $(5)$ Time $2^{24}$ [10]
+CPA using Multi-bit Selection Function Full 22-bit selection function =7200=7200 traces $3$-bit selection function =8200=8200 traces [47]
+Differential Fault Attack Full Fault queries $2^{10}$ [48]
+Collision Attack 22 Time $2^{98}$ [49]
+Free-start Collison Attack 33 Time $2^{10}$ [50]
+GIFT-128 / COFB MILP-Based Linear Attack 2020 $65$-bit key recovery [18]
+Differential Key Recovery Attack $23$ Time $2^{120}$, Data $2^{86}$ [51]
+Integral Distinguisher using Division Property $11$ - [20]
+Differential Attack $26$ Time $2^{123.245}$ [21]
+Photon-Beetle Key Recovery Irregularity Full Data $2^{50}$ bytes [28]
+Differential Fault Attack Full Time $2^{16}$, Fault queries $2^{37.15}$ [52]
+Committing Attack Initial State No queries [53]
+Pre-image Attack $3.5$ Time = $2^{112}$, Data $2^{65}$ [54]
+Elephant Interpolation Attack 88 Time $2^{70}$ [33]
+Fault Analysis (Dumbo) Full 8525085-250 ciphertexts [34]
+Blind SCA on LFSR Full <2<2 days (practical) [35]
+Romulus Impossible Differential $22$ Time $2^{373.48}$ , Data $2^{92.22}$ [38]
+Matching / Authenticity Attack Full - [39]
+Meet-in-the-Middle (MITM) $23$ Time $2^{376}$ , Data $2^{104}$ [40]
+Impossible Boomerang $27$ $28$ Time $2^{337}$, Data $2^{131.3}$ Time $2^{382.8}$, Data $2^{130.26}$ [55]
+Differential MITM 25 Time $2^{366}$, Data $2^{122.3}$ [56]
+Impossible Differential MITM 2121 $22$ $23$ $27$ $28$ Time $2^{344.33}$, Data $2^{122.89}$ Time $2^{378.22}$, Data $2^{121.50}$ Time $2^{378.22}$, Data $2^{121.50}$ Time = $2^{361.06}$, Data = $2^{124.99}$ Time $2^{378.22}$, Data $2^{129.50}$ [57]
+Shannon’s property confusion in any symmetric cipher is mainly dependent on the non-linear component used.
+The most commonly used component to achieve non-linearity in symmetric cryptography is a substitution box (S-box).
+Analysing the cryptographic profile of an S-box plays a vital role in determining the overall strength of a cryptosystem.
+The literature reports nearly $24$ cryptographic properties [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] that can be used as a tool to estimate the quality of an S box.
+Some of the attacks mentioned earlier exploit specific S-box properties in order to devise practical attack characteristics for a particular cipher.
+This paper aims to provide a complete picture of the cryptographic properties of S-boxes, their relevance to known attacks, and related security limits  [^1] Current paper is an extended version of our previous paper [74].
+Our primary goal is to help researchers better understand and evaluate the security profile of S-boxes in a principle manner.
 Section 2 of this paper elaborates the specifications of six finalists of NIST LWC.
-Section 3 discusses the fundamental properties of sboxes and classifies them on the basis of relevant attack category.
-Secondly, upper and lower bounds on these properties are highlighted along with their behaviour under affine equivalence.
-In Section 4 the results of sbox analysis are discussed in detail.
-Conclusion is presented in Section 5.
+Section 3 discusses the fundamental properties of S boxes and classifies them according to the relevant attack category.
+Secondly, the upper and lower bounds of these properties are highlighted along with their behavior under affine equivalence.
+In Section 4, the results of the S-box analysis are discussed in detail.
+In Section 5, we generate a new 44-bit S-box with optimal cryptographic properties using the PEIGEN S-box generation tool [75], and draw a comparison of the properties of this S-box with the LWC finalists having 44-bit S-boxes.
+Section 6 discusses and highlights the scope and limitations of this paper.
+Finally, we conclude our findings in Section 7.
 2 Overview of Finalists
-This section provides a brief introduction to each of the $6$ finalists from NIST lightweight standardization that make use of sbox as one of the design components.
+This section provides a brief introduction to each of the $6$ finalists from NIST lightweight standardization that make use of S-box as one of the design components.
 2.1 ASCON
@@
 [76].
-ASCON comprises of three AEAD variants: $128$, $128$ ($a$ and $pq$), and two hashing algorithms: Hash and Hash-$a$.
-Its main functionality comes from the use of a $320$ bit permutation which is combined with different constants and number of rounds, depending upon the variant.
-In addition, to achieve non-linearity in the encryption algorithm a $5\times 5$ affine equivalent sbox of $\chi$ mapping from Keccak permutation is used.
+ASCON comprises of three authenticated encryption and associated data (AEAD) variants: $128$, $128$ ($a$ and $pq$), and two hashing algorithms: Hash and Hash-$a$.
+Its main functionality comes from the use of a $320$ bit permutation, which is combined with different constants and number of rounds, depending upon the variant.
+In addition, to achieve non-linearity in the encryption algorithm a $5\times 5$ affine equivalent S-box of $\chi$ mapping from Keccak permutation is used.
 The ASCON family has been selected by NIST for lightweight standardization because of its security strength and flexibility to be used in multiple applications while bearing a minimum implementation cost [6].
@@
 [77], it is a nonce based AEAD cryptographic scheme designed to resist against implementation based attacks at an optimized implementation cost.
-ISAP uses $400$ bit Keccak and $320$ bit ASCON permutations in its design alongside $5\times 5$ bit ASCON sbox as a non-linear component.
+ISAP uses $400$ bit Keccak and $320$ bit ASCON permutations in its design alongside $5\times 5$ bit ASCON S-box as a non-linear component.
 2.3 GIFT-COFB
@@
 Whereas the output is a $128$ bit tag and ciphertext of same length as the input message.
-A $4\times 4$ bit sbox is used as the confusion component.
+A $4\times 4$ bit S-box is used as the confusion component.
 2.4 Photon-Beetle
 Photon-Beetle is an AEAD based hashing and encryption scheme presented by J.
-Guo et al.[65].
+Guo et al.
+[79].
 The encryption routine comprises of $12$ rounds consisting of $4$ separate stages i.e.
 XOR constant, substitute cells, shift rows and serial mix columns.
-Photon-Beetle (AEAD) uses the $4$ $4$ bit sbox of Present block cipher in order to achieve non-linearity in its design.
+Photon-Beetle (AEAD) uses the $4$ × $4$ bit S-box of Present block cipher in order to achieve non-linearity in its design.
 2.5 Elephant
@@
 These encryption variants take $128$ bit key and $96$ bit nonce, however the block sizes are $160$, $176$, $200$.
-A $4\times 4$ sbox as the non-linear part.
-In the $80$ rounds of encryption there are $160$ differentially active sboxes.
+A $4\times 4$ S-box as the non-linear part.
+In the $80$ rounds of encryption there are $160$ differentially active S-boxes.
 2.6 Romulus
@@
 The encryption algorithm takes $128$ bits of key and nonce, and runs $40$ rounds of encryption.
-Skinny-$128$ tweakable block cipher [68] uses $8\times 8$ sbox to achieve non-linearity in its design.
-Table 2: Classification of s-box properties as per
-relevance to existing cryptanalysis techniques
+Skinny-$128$ tweakable block cipher [82] uses $8\times 8$ S-box to achieve non-linearity in its design.
+Table 3: Classification of S-box properties as per relevance to existing cryptanalysis techniques
 S# Criteria Properties
-1 Generic
+1 Generic Balancedness; Permutation; Order of Permutation; Fixed Points (FP); Opposite Fixed Points (OFP); Bit Independence Criteria (BIC); Strict Avalanche Criteria (SAC); Auto Correlation Table (ACT) [83]; Absolute Indicator (AI) [84]; Sum of Square Indicators (SSI) [84]
+2 Linear Cryptanalysis Nonlinearity (NL) [62, 63]; Linear Approximation Table (LAT) [64, 85]; Linear Approximation Probability (LAP); Linear Branch Number (LBN); Linear Structures (LS) [65, 66]; Correlation Immunity (CI) [86]
+3 Differential Cryptanalysis Differential Distribution Table (DDT) [64]; Differential Uniformity (DU) [62, 63]; Differential Branch Number (DBN); Propagation Characteristics (PC) [87]; Undisturbed Bits [68]
+4 Boomerang Cryptanalysis Boomerang Connectivity Table (BCT); Boomerang Uniformity (BU) [67]
+5 Differential–Linear Cryptanalysis Linear Structures (LS) [65, 66]; Differential Linear Connectivity Table (DLCT) [69]; Differential Linear Uniformity (DLU) [70]
+6 Algebraic Cryptanalysis Algebraic Degree (AD) [71]; Interpolation Polynomial (IP)
+7 Side-Channel Analysis DPA-SNR [72]; Transparency Order (TO) [73]
+Table 4: Theoretical Bounds on the Cryptographic Properties of S-boxes
+S# Properties Bounds
+For
+4×4
+For
+5×5
+For
+8×8
+Ideal
+Value
+1 $\mathsf{OP}$ $1\leq\mathsf{OP}\leq 2^{n}$ $1\leq\mathsf{OP}\leq 16$ $1\leq\mathsf{OP}\leq 32$ $1\leq\mathsf{OP}\leq 256$ Close to UB
+2 $\mathsf{FP}$ $0\leq\mathsf{FP}\leq 2^{n}$ $0\leq\mathsf{FP}\leq 16$ $0\leq\mathsf{FP}\leq 32$ $0\leq\mathsf{FP}\leq 256$ Close to LB
+3 $\mathsf{OFP}$ $0\leq\mathsf{OFP}\leq 2^{n}$ $0\leq\mathsf{OFP}\leq 16$ $0\leq\mathsf{OFP}\leq 32$ $0\leq\mathsf{OFP}\leq 256$ Close to LB
+4 $\mathsf{BIC}$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ 0
+5 $\mathsf{SAC}$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0.5$
+6 $\barwedge_{S}$ $\sqrt{\tfrac{2^{2n}}{2^{n}-1}}\leq\barwedge_{S}\leq 2^{n}$ $0<\barwedge_{S}\leq 16$ $0<\barwedge_{S}\leq 32$ $0<\barwedge_{S}\leq 256$ Close to LB
+7 $\doublebarwedge_{S}$ $2^{2n}\leq\doublebarwedge_{S}\leq 2^{3n+m}$ $256\leq\doublebarwedge_{S}\leq 65536$ $1024\leq\doublebarwedge_{S}\leq 1048576$ $65536\leq\doublebarwedge_{S}\leq 4294967296$ Close to LB
+8 $\mathsf{LAT}$ $0\leq\mathsf{LAT}\leq 2^{n}-1$ $0\leq\mathsf{LAT}\leq 15$ $0\leq\mathsf{LAT}\leq 31$ $0\leq\mathsf{LAT}\leq 255$ -
+9 $\xi$ $0\leq\xi\leq\tfrac{1}{2}$ $0\leq\xi\leq\tfrac{1}{2}$ $0\leq\xi\leq\tfrac{1}{2}$ $0\leq\xi\leq\tfrac{1}{2}$ Close to LB
+10 $\mathsf{NL}$
+For $n$ even: $0\leq\mathsf{NL}<2^{n-1}-2^{\tfrac{n}{2}-1}$
+For $n$ odd: $0\leq\mathsf{NL}<2^{n-1}-2^{\tfrac{n-1}{2}}$
+$0\leq\mathsf{NL}<6$ $0\leq\mathsf{NL}<12$ $0\leq\mathsf{NL}<120$ Close to UB
+11 $\mathsf{BN_{L}}$ $2\leq\mathsf{BN_{L}}\leq n-1$ $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{L}}\leq 4$ $2\leq\mathsf{BN_{L}}\leq 7$ Close to UB
+12 $\mathsf{LS}$ $0\leq\mathsf{LS}\leq 2^{n+m}$ $0\leq\mathsf{LS}\leq 256$ $0\leq\mathsf{LS}\leq 1024$ $0\leq\mathsf{LS}\leq 65536$ Close to LB
+13 $\mathsf{CI}$ $0\leq\mathsf{CI}\leq n$ $0\leq\mathsf{CI}\leq 4$ $0\leq\mathsf{CI}\leq 5$ $0\leq\mathsf{CI}\leq 8$ Close to LB
+14 $\mathsf{DDT}$ $0\leq\mathsf{DDT}\leq 2^{n}-1$ $0\leq\mathsf{DDT}\leq 15$ $0\leq\mathsf{DDT}\leq 31$ $0\leq\mathsf{DDT}\leq 255$ -
+15 $\mathsf{DU}$ $2\leq\mathsf{DU}\leq 2^{n}$ $2\leq\mathsf{DU}\leq 16$ $2\leq\mathsf{DU}\leq 32$ $2\leq\mathsf{DU}\leq 256$ Close to LB
+16 $\mathsf{BN_{D}}$ $2\leq\mathsf{BN_{D}}\leq\lceil\tfrac{2n}{3}\rceil$ $2\leq\mathsf{BN_{D}}\leq 3$ $2\leq\mathsf{BN_{D}}\leq 4$ $2\leq\mathsf{BN_{D}}\leq 6$ Close to UB
+17 $\mathsf{PC}$ $0\leq\mathsf{PC}\leq n$ $0\leq\mathsf{PC}\leq 4$ $0\leq\mathsf{PC}\leq 5$ $0\leq\mathsf{PC}\leq 8$ Close to UB
+18 $\mathsf{UDB}$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ 0
+19 $\beta\mathsf{U}$ $2\leq\beta\mathsf{U}\leq 2^{n}$ $2\leq\beta\mathsf{U}\leq 16$ $2\leq\beta\mathsf{U}\leq 32$ $2\leq\beta\mathsf{U}\leq 256$ Close to LB
+20 $\mathsf{DLU}$ $2^{\tfrac{n}{2}-1}<\mathsf{DLU}\leq 2^{n-1}$ $2<\mathsf{DLU}\leq 8$ $4<\mathsf{DLU}\leq 16$ $8<\mathsf{DLU}\leq 128$ Close to LB
+21 $\mathsf{AD}$ $1\leq\mathsf{AD}\leq n-1$ $1\leq\mathsf{AD}\leq 3$ $1\leq\mathsf{AD}\leq 4$ $1\leq\mathsf{AD}\leq 7$ Close to UB
+22 $\mathsf{DPA_{SNR}}$
+For balanced S-box:
+$1\leq\mathsf{DPA_{SNR}}\leq 2^{\tfrac{n}{2}}$
+$0\leq\mathsf{DPA_{SNR}}\leq 4$ $0\leq\mathsf{DPA_{SNR}}<6$ $0\leq\mathsf{DPA_{SNR}}\leq 16$ Close to LB
+23 $\mathsf{TO}$ $0\leq\mathsf{TO}\leq m$ $0\leq\mathsf{TO}\leq 4$ $0\leq\mathsf{TO}\leq 5$ $0\leq\mathsf{TO}\leq 8$ Close to LB
+Table 5: S-box Properties under Affine Equivalence
+Property Affine Variant Affine Invariant
+Bijectivity
 Balancedness
 Permutation
-Order of Permutation
-Fixed Points (FP)
-Opposite Fixed Points (OFP)
-Bit Independence Criteria (BIC)
-Strict Avalanche Criteria (SAC)
-Auto Correlation Table (ACT) [69]
-Absolute Indicator (AI) [70]
-Sum of Square Indicators (SSI) [70]
-2 Linear Cryptanalysis
-Nonlinearity (NL) [50, 51]
-Linear Approximation Table (LAT) [52, 71]
-Linear Approximation Probability (LAP)
-Linear Branch Number (LBN)
-Linear Structures (LS) [53, 54]
-Correlation Immunity (CI) [72]
-3 Differential Cryptanalysis
-Differential Distribution Table (DDT) [52]
-Differential Uniformity (DU) [50, 51]
-Differential Branch Number (DBN)
-Propagation Characteristics (PC) [73]
-Undisturbed Bits [56]
-4 Boomerang Cryptanalysis
-Boomerang Connectivity Table (BCT)
-Boomerang Uniformity (BU) [55]
-4 Differential - Linear Cryptanalysis
-Linear Structures (LS) [53, 54]
-Differential Linear Connectivity Table (DLCT) [57]
-Differential Linear Uniformity (DLU) [58]
-5 Algebraic Cryptanalysis
-Algebraic Degree (AD) [59]
-Interpolation Polynomial (IP)
-6 Side Channel Analysis
-DPA-SNR[60]
-Transparency Order (TO) [61]
-Table 3: Theoretical Bounds on the Cryptographic Properties of Sboxes
+$\mathsf{OP}$
+$\mathsf{FP}$
+$\mathsf{OFP}$
+$\mathsf{BIC}$
+$\mathsf{SAC}$
+$\mathsf{ACT}$
+[…]
+[…]
+$\mathsf{NL}$
+$\mathsf{LAT}$
+$\mathsf{BN\_L}$
+$\mathsf{LS}$
+$\mathsf{CI}$
+$\mathsf{DDT}$
+$\mathsf{DU}$
+$\mathsf{BN\_D}$
+$\mathsf{PC}$
+$\mathsf{UDB}$
+$\mathsf{BCT}$
+$\beta\mathsf{U}$
+$\mathsf{DLCT}$
+$\mathsf{DLU}$
+$\mathsf{AD}$
+$\mathsf{IP}$
+$\mathsf{DPA\text{-}SNR}$
+$\mathsf{TO}$
+[…]
+Table 6: Results of S-box Analysis – $8\times 8$ and $5\times 5$ bit case
   S# Properties
-Bounds
-For
-$4\times 4$ Sbox
-For
-$5\times 5$ Sbox
-For
-$8\times 8$ Sbox
-Ideal Value
-1
-$\mathsf{OP}$
-$1\leq\mathsf{OP}\leq 2^{n}$ $1\leq\mathsf{OP}\leq 16$ $1\leq\mathsf{OP}\leq 32$ $1\leq\mathsf{OP}\leq 256$ Close to LB
-2
-$\mathsf{FP}$
-$0\leq\mathsf{FP}\leq 2^{n}$ $0\leq\mathsf{FP}\leq 16$ $0\leq\mathsf{FP}\leq 32$ $0\leq\mathsf{FP}\leq 256$ Close to LB
-3
-$\mathsf{OFP}$
-$0\leq\mathsf{OFP}\leq 2^{n}$ $0\leq\mathsf{OFP}\leq 16$ $0\leq\mathsf{OFP}\leq 32$ $0\leq\mathsf{OFP}\leq 256$ Close to LB
-4
-$\mathsf{BIC}$
-$0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ 00
-5
-$\mathsf{SAC}$
-$0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0.5$
-6
-$\barwedge_{S}$
-$\sqrt{{\frac{2^{2n}}{2^{n}-1}}}\leq\barwedge_{S}\leq 2^{n}$
-$0<\barwedge_{S}\leq 16$
-$0<\barwedge_{S}\leq 32$
-$0<\barwedge_{S}\leq 256$
-Close to LB
-7
-$\doublebarwedge_{S}$
-$2^{2n}\leq\doublebarwedge_{S}\leq 2^{(3n+m)}$ $256\leq\doublebarwedge_{S}\leq 65536$ $1024\leq\doublebarwedge_{S}\leq 1048576$ $65536\leq\doublebarwedge_{S}\leq 4294967296$ Close to LB
-8
-$\mathsf{LAT}$
-$0\leq\mathsf{LAT}\leq 2^{n-1}$ $0\leq\mathsf{LAT}\leq 8$ $0\leq\mathsf{LAT}\leq 16$ $0\leq\mathsf{LAT}\leq 128$ Close to LB
-9
-$\xi$
-$0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ Close to LB
-10
-$\mathsf{NL}$
-For $n:even$
-$0\leq\mathsf{NL}<2^{n-1}-2^{\frac{n}{2}-1}$
-For $n:odd$
-$0\leq\mathsf{NL}<2^{n-1}-2^{\frac{n-1}{2}}$
-$60\leq\mathsf{NL}<$ 6 $120\leq\mathsf{NL}<$ 12 $1200\leq\mathsf{NL}<$ 120 Close to UB
-11
-$\mathsf{BN_{L}}$
-$12\leq\mathsf{BN_{L}}\leq n-$ 1 $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{L}}\leq 4$ $2\leq\mathsf{BN_{L}}\leq 7$ Close to UB
-12
-$\mathsf{LS}$
-$0\leq\mathsf{LS}\leq 2^{n+m}$ $0\leq\mathsf{LS}\leq 256$ $0\leq\mathsf{LS}\leq 1024$ $0\leq\mathsf{LS}\leq 65536$ Close to LB
-13
-$\mathsf{CI}$
-$0\leq\mathsf{CI}\leq n$ $0\leq\mathsf{CI}\leq 4$ $0\leq\mathsf{CI}\leq 5$ $0\leq\mathsf{CI}\leq 8$ Close to LB
-14
-$\mathsf{DU}$
-$2\leq\mathsf{DU}\leq 2^{n}$ $2\leq\mathsf{DU}\leq 16$ $2\leq\mathsf{DU}\leq 32$ $2\leq\mathsf{DU}\leq 256$ Close to LB
-15
-$\mathsf{BN_{D}}$
-$2\leq\mathsf{BN_{D}}\leq\lceil{\frac{2n}{3}}\rceil$ $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{D}}\leq 4$ $2\leq\mathsf{BN_{D}}\leq 6$ Close to UB
-16
-$\mathsf{PC}$
-$0\leq\mathsf{PC}\leq n$  $0\leq\mathsf{PC}\leq 4$ $0\leq\mathsf{PC}\leq 5$ $0\leq\mathsf{PC}\leq 8$ Close to UB
-17
-$\mathsf{UDB}$
-$\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ 00
-18
-$\beta\mathsf{U}$
-$2\leq\beta\mathsf{U}\leq 2^{n}$ $2\leq\beta\mathsf{U}\leq 16$ $2\leq\beta\mathsf{U}\leq 32$ $2\leq\beta\mathsf{U}\leq 256$ Close to LB
-19
-$\mathsf{DLU}$
-$2^{\frac{n}{2}-1}<\mathsf{DLU}\leq 2^{n-1}$ $2<\mathsf{DLU}\leq 8$ $4<\mathsf{DLU}\leq 16$ $8<\mathsf{DLU}\leq 128$ Close to LB
-20
-$\mathsf{AD}$
-$11\leq\mathsf{AD}\leq n-$ 1 $1\leq\mathsf{AD}\leq 3$ $1\leq\mathsf{AD}\leq 4$ $1\leq\mathsf{AD}\leq 7$ Close to UB
-21
-$\mathsf{DPA_{SNR}}$
-For balanced S-box:
-$1\leq\mathsf{DPA_{SNR}}\leq 2^{\frac{n}{2}}$
-$0\leq\mathsf{DPA_{SNR}}\leq 4$ $60\leq\mathsf{DPA_{SNR}}<$ 6 $0\leq\mathsf{DPA_{SNR}}\leq 16$ Close to LB
-22
-$\mathsf{TO}$
-$0\leq\mathsf{TO}\leq m$ $0\leq\mathsf{TO}\leq 4$ $0\leq\mathsf{TO}\leq 5$ $0\leq\mathsf{TO}\leq 8$ Close to LB
-3 Cryptographic Properties of Sboxes and their Classification
-The strength of cryptographic primitives is best analysed with the help of well known cryptanalysis techniques.
+Romulus
+$8\times 8$ Skinny-128
+ASCON/ISAP
+$5\times 5$
+  1 Bijectivity
+  2 Balancedness
+  3 Permutation
+  4 $\mathsf{OP}$ 140 26
+  5 $\mathsf{FP}$ 1 0
+  6 $\mathsf{OFP}$ 0 0
+  7 $\mathsf{BIC}$ 1.000000 1.000000
+  8 $\mathsf{SAC}$ 1.000000 1.000000
+  9 $\barwedge_{S}$ 256 32
+  10 $\doublebarwedge_{S}$ 4194304 8192
+  11 $\xi$ 0.25000 0.25000
+  12 $\mathsf{NL}$ 64 8
+  13 $\mathsf{BN_{L}}$ 2 3
+  14 $\mathsf{LS}$ 601 91
+  15 $\mathsf{CI}$ 0 0
+  16 $\mathsf{DU}$ 64 8
+  17 $\mathsf{BN_{D}}$ 2 3
+  18 $\mathsf{PC}$ 0 0
+  19 $\mathsf{UDB}$ 258 35
+  20 $\beta\mathsf{U}$ 256 16
+  21 $\mathsf{DLU}$ 128 16
+  22 $\mathsf{AD}$ 6 2
+  23 $\mathsf{DPA\mbox{-}SNR}$ 6.312455 3.015113
+  24 $\mathsf{TO}$ 7.174510 4.258065
+Table 7: Results of S-box Analysis – $4\times 4$ bit case
+  S# Properties
+New Sbox
+$4\times 4$
+Elephant
+$4\times 4$
+GIFT-COFB
+$4\times 4$ GIFT
+Photon-Beetle
+$4\times 4$ Present
+  1 Bijectivity
+  2 Balancedness
+  3 Permutation
+  4 $\mathsf{OP}$ 7 13 9 7
+  5 $\mathsf{FP}$ 3 0 0 0
+  6 $\mathsf{OFP}$ 1 1 1 1
+  7 $\mathsf{BIC}$ 0.577350 1.000000 1.000000 1.000000
+  8 $\mathsf{SAC}$ 0.7500 1.000000 1.000000 1.000000
+  9 $\barwedge_{S}$ 8 16 16 16
+  10 $\doublebarwedge_{S}$ 640 1024 1024 1024
+  11 $\xi$ 0.25000 0.25000 0.25000 0.25000
+  12 $\mathsf{NL}$ 4 4 4 4
+  13 $\mathsf{BN_{L}}$ 2 2 2 2
+  14 $\mathsf{LS}$ 0 9 9 9
+  15 $\mathsf{CI}$ 0 0 0 0
+  16 $\mathsf{DU}$ 4 4 6 4
+  17 $\mathsf{BN_{D}}$ 2 3 2 3
+  18 $\mathsf{PC}$ 0 0 0 0
+  19 $\mathsf{UDB}$ 0 3 6 6
+  20 $\beta\mathsf{U}$ 10 16 16 16
+  21 $\mathsf{DLU}$ 4 8 8 8
+  22 $\mathsf{AD}$ 3 3 3 3
+  23 $\mathsf{DPA\mbox{-}SNR}$ 2.578634 2.398501 2.398501 2.128608
+  24 $\mathsf{TO}$ 3.400000 3.266667 3.466667 3.533333
+3 Cryptographic Properties of S-boxes and their Classification
+The strength of cryptographic primitives is best analysed with the help of well-known cryptanalysis techniques.
 Shannon’s theory of confusion and diffusion plays the foundational role in the design and analysis of iterative ciphers.
-The aspect of confusion is introduced in symmetric ciphers using highly non-linear components known as sboxes.
+The aspect of confusion is introduced in symmetric ciphers using highly non-linear components such as S-boxes.
 Designers and cryptanalysts treat these ciphers as a white box, analysing all the underlying mathematical components in order to understand and rule out any potential attack vectors, by imposing design practices based on standard characteristics and theoretically well-defined bounds.
-Owing to the importance of sboxes in the design of cryptographic algorithms, an extensive research has been done over the years to explain the most refined design and analysis criteria that could be employed in order to secure them against potential attack vectors.
+Owing to the importance of S-boxes in the design of cryptographic algorithms, an extensive research has been done over the years to explain the most refined design and analysis criteria that could be employed in order to secure them against potential attack vectors.
 These properties can be classified in a number of ways depending whether upon the types of representation in boolean algebra or the statistical and implementation based attacks that could be launched through exploitation of relevant metrics or via studying the inter-dependencies that link one property to another.
-This section briefly covers an introduction to around twenty-five cryptographic properties of sboxes that hold significance in the current literature.
+This section briefly covers an introduction to around twenty-four cryptographic properties of S-boxes that hold significance in the current literature.
 Each of the properties is related to one or more crytanalysis techniques, that could exploit the non-linear part of a block cipher, see Table 3.
-In addition, Table 3 presents the upper and lower bounds (UB & LB) on these properties with respect to the dimension of sbox being used [74, 75, 59, 76, 77, 78, 60, 61].
+In addition, Table 4 presents the upper and lower bounds (UB & LB) on these properties with respect to the dimension of S-box being used [88, 89, 71, 90, 91, 92, 72, 73].
 We also talk about the behaviour of these properties under affine equivalence, see Table 5.
-The properties along with their respective bounds and equivalence characteristics aid designers and cryptanalysts in the selection of good sboxes.
+The properties along with their respective bounds and equivalence characteristics aid designers and cryptanalysts in the selection of good S-boxes.
 3.1 Diffusion Characteristics
-Properties such as balancedness, bijectivity, order of permutation, fixed and opposite fixed points, strict avalanche and bit independence criteria, absolute and sum of square indicator are general pre-requisites that should adhere to defined bounds for any sbox in order to avoid correlations between the input and output bits and serve as a pre-requisite to ensure security against various statistical cryptanalysis attacks.
+Properties such as balancedness, bijectivity, order of permutation, fixed and opposite fixed points, strict avalanche and bit independence criteria, absolute and sum of square indicator are general pre-requisites that should adhere to defined bounds for any S-box in order to avoid correlations between the input and output bits and serve as a pre-requisite to ensure security against various statistical cryptanalysis attacks.
 Balancedness
-A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is balanced if it has an equal number of $1$’s and 00’s in its coordinate and component functions. Mathematically;
+A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is balanced if it has an equal number of $1$’s and 0’s in its coordinate and component functions. Mathematically;
 $P(0)=P(1)=\frac{1}{2}=50\%$ (1)
 An imbalanced boolean mapping stays at a risk of achieving high linear approximation probabilities [93].
-Balancedness is invariant when under affine equivalence, which means that the number of 00’s ad $1$’s is preserved in $S$ if an affine transformation is applied to it [80, 81, 82].
+Balancedness is invariant when under affine equivalence, which means that the number of 0’s ad $1$’s is preserved in $S$ if an affine transformation is applied to it [94, 95, 96].
 Permutation
@@
 Bit Independence Criteria (BIC)
-A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is said to satisfy BIC, if inverting an input bit $u$ has a changing effect on the output bit $v$ i.e.
-all the avalanche variables $(\gamma_{u},\gamma_{v},....)$ representing difference between bits of input and output of $S$ become pairwise independent [85].
-Mathematically, BIC is computed using correlation co-efficient of the avalanche variables:-
+[…]
 $\rho(\gamma_{u},\gamma_{v})=\frac{cov(\gamma_{u},\gamma_{v})}{\sigma(\gamma_{u})\sigma(\gamma_{v})}$ (4)
-[…]
-Table 4: S-box Properties under Affine Equivalence
-Property Affine Variant Affine Invariant
-Bijectivity
-Balancedness
-Permutation
-$\mathsf{OP}$
-$\mathsf{FP}$
-$\mathsf{OFP}$
-$\mathsf{BIC}$
-$\mathsf{SAC}$
-$\mathsf{ACT}$
-[…]
-[…]
-$\mathsf{NL}$
-$\mathsf{LAT}$
-$\mathsf{BN_{L}}$
-$\mathsf{LS}$
-$\mathsf{CI}$
-$\mathsf{DDT}$
-$\mathsf{DU}$
-$\mathsf{BN_{D}}$
-$\mathsf{PC}$
-$\mathsf{UDB}$
-$\mathsf{BCT}$
-$\beta\mathsf{U}$
-$\mathsf{DLCT}$
-$\mathsf{DLU}$
-$\mathsf{AD}$
-$\mathsf{IP}$
-$\mathsf{DPA-SNR}$
-$\mathsf{TO}$
-For the coordinate and component functions of $S$ the BIC value should stay in the range of $[0,1]$.
-An ideal sbox $S$ has a BIC value equal to 00.
+Where; $u\neq v\$and u,v=1,2,..,m\ u,v=1,2,.....,m\ .
+The numerator in eq 4 denotes co-variance between two avalanche variables ($\gamma_{u}$ and $\gamma_{v}$), while $\sigma(\gamma_{u})$ and $\sigma(\gamma_{u})$ in the denominator is the product of the same variables.
+For the coordinate and component functions of $S$, the BIC value should stay in the range of $[0,1]$.
+An ideal S-box has a BIC value equal to 0.
 If an affine transformation is applied to $S$, BIC does not stay invariant [100].
@@
 $S_{\gamma}=S(\varkappa)\oplus S(\varkappa\oplus\gamma)$ (5)
-Where $\gamma\in F_{2}^{n},\gamma\neq 0,wt(\gamma)=1$ SAC is variant under affine transformation, which means that a function not satisfying SAC could be transformed to a function that satisfies SAC using affine equivalence [87].
+Where $\gamma\in F_{2}^{n},\gamma\neq 0,wt(\gamma)=1$.
+SAC is variant under affine transformation, which means that a function not satisfying SAC could be transformed to a function that satisfies SAC using affine equivalence [101].
 Auto Correlation Transform (ACT)
@@
 The overall impact of auto correlation of a boolean function or a vectorial boolean function is studied by finding the maximum value in ACT spectrum, known as absolute indicator (AI).
-Let $f$ be a Boolean function over $F_{2}^{n}$ i.e. $f:F_{2}^{n}\rightarrow F_{2}$. The AI of $f$ is defined as [74]:-
+Let $f$ be a Boolean function over $F_{2}^{n}$ i.e.
+$f:F_{2}^{n}\rightarrow F_{2}$.
+The AI of $f$ is defined as [88]:-
 $\barwedge_{f}=\max_{\gamma\in F_{2}^{n}\backslash\{0\}}|\triangle_{f}(\gamma)|$ (7)
@@
 $F(D_{\gamma}f)=\triangle_{f}(\gamma)=\sum_{\varkappa\in F_{2}^{n}}(-1)^{D_{\gamma}f}$ (8)
-Here $D_{\gamma}f=f(x)\oplus f(x\oplus a)$.
+Here $D_{\gamma}f=f(\varkappa)\oplus f(\varkappa\oplus a)$.
 For a vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$, using (6) the AI of $S$ is defined as [102]:-
@@
 $\barwedge_{S}$ stays invariant under affine transformation.
-For balanced vectorial boolean functions, the value of AI is always 00.
+For balanced vectorial boolean functions, the value of AI is always 0.
 If a boolean mapping has high $\barwedge_{S}$, it becomes highly susceptible to cube attacks [103, 88].
@@
 $\doublebarwedge_{S}=\sum_{\gamma\in F_{2}^{n},,\rho\in F_{2}^{m}}\triangle^{2}_{S}(\gamma,\rho)$ (10)
-SSI is not variant under affine equivalence [90]. For achieving good diffusion characteristics in a cryptographic algorithm, sboxes should have low SSI and AI values [91, 74].
+SSI is not variant under affine equivalence [104]. For achieving good diffusion characteristics in a cryptographic algorithm, S-boxes should have low SSI and AI values [105, 88].
 3.2 Linear Cryptanalysis
@@
 Block ciphers having a highly non-linear layer are resistant towards linear cryptanalysis.
-To study the number of active sboxes in the linear trails of a block cipher, J.
+To study the number of active S-boxes in the linear trails of a block cipher, J.
 Daemon et al.
 [107] introduced linear branch number.
-Theoretically, in order to resist linear (and its variant) cryptanalysis techniques an sbox should have a low linear approximation probability, high non-linearity, high order of linear branch number, equal to none linear structures, and low order correlation immunity.
+Theoretically, in order to resist linear (and its variant) cryptanalysis techniques an S-box should have a low linear approximation probability, high non-linearity, high order of linear branch number, equal to none linear structures, and low order correlation immunity.
 Linear Approximation Table (LAT)
@@
 LAT can be mathematically represented as follows:-
-$\mathsf{LAT}(\varrho,\psi)=\{\#\varkappa\ |\ \varrho\cdot\varkappa=\psi\cdot S(\varkappa)\}$ (11)
-Definition 3.1
+$\mathsf{LAT}(\varrho,\psi)=|\{\#\varkappa\ |\ \varrho\cdot\varkappa=\psi\cdot S(\varkappa)\}|-2^{n-1}$ (11)
+Definition 1.
 The maximum LAT table entry is measured in the form of Linear Approximation Probability (LAP), mathematically given as follows:-
-$\xi=\frac{1}{2^{n}}\max_{\varrho,\psi\in F_{2}^{n},F_{2}^{m}\backslash\{0\}}\bigg{|}\sum_{\varkappa=1}^{2^{n-1}}\mathsf{LAT}(\varrho,\psi)\bigg{|}$ (12)
+$\xi=\frac{1}{2^{n}}\max_{\varrho,\psi\in F_{2}^{n},F_{2}^{m}\backslash\{0\}}\bigg|\mathsf{LAT}(\varrho,\psi)\bigg|$ (12)
 The LAT table is of size $2^{n}\times 2^{m}$.
@@
 Linear Branch Number (LBN)
-In block cipher design strategy, branch number is a significant property used to measure the strength of an sbox against linear cryptanalysis.
-An sbox with high branch number could increase the diffusion power of a cryptographic algorithm.
+In block cipher design strategy, branch number is a significant property used to measure the strength of an S-box against linear cryptanalysis.
+An S-box with high branch number could increase the diffusion power of a cryptographic algorithm.
 Mathematically:-
 $\mathsf{BN_{L}(S)}=\min_{\gamma\in F_{2}^{n},\rho\in F_{2}^{m}\backslash\{0\}}\{w_{t}(\gamma)+w_{t}(\rho)\}$ (16)
-LBN is variant under affine transformation i.e. two sboxes can be affine equivalent of each other but they can have different values of LBN [98].
+LBN is variant under affine transformation i.e. two S-boxes can be affine equivalent of each other but they can have different values of LBN [112].
 Linear Structures (LS)
@@
 A block cipher has Linear Structures (LS) when there exist subsets of input, output and key bits, such that changing all the input and key bits has the same effect on the XOR of the output bits.
-Definition 3.2
+Definition 2.
 Let $f:F_{2}^{n}\rightarrow F_{2}$ be a Boolean function, then for some $\gamma\in F_{2}^{n},f(\varkappa\oplus\gamma)\oplus f(\varkappa)=c$ holds for all $\varkappa\in F_{2}^{n}$ where $c\in F_{2}$ is a constant, then $\gamma$ is called a LS of $f(\varkappa)$.
 More specifically, $\gamma$ is called an invariant LS of $f(\varkappa)$ if $c=0$ and a complementary LS of $f(\varkappa)$ if $c=1$ [105].
-Definition 3.3
+Definition 3.
 Let $S:F_{2}^{n}\rightarrow F_{2}^{m}$ be a vectorial boolean mapping if for some $\gamma\in F_{2}^{n},S(\varkappa\oplus\gamma)\oplus S(\varkappa)=c$ holds for all $\varkappa\in F_{2}^{n}$, where $c\in F_{2}^{m}$ is a constant, then $\gamma$ is called a LS of $S(\varkappa)$.
 More specifically, $\gamma$ is called an invariant LS of $S(\varkappa)$ if $c=0$ and a complementary LS of $S(\varkappa)$ if $c=1$ [113].
-Sboxes having LS are vulnerable to attacks more robust than exhaustive key search, such as differential and linear cryptanalysis [100].
+S-boxes having LS are vulnerable to attacks more robust than exhaustive key search, such as differential and linear cryptanalysis [114].
 LS is an affine variant property of any boolean mapping $S$ [115, 88].
@@
 So using (14), we can say that $S$ is $k$-order CI when [116]:-
-[…]
-CI shows variant behaviour when an affine transformation is applied to $S$ [74].
-For a good cryptographic sbox the lower is the order of CI the better.
+$\{\mathsf{W}(\gamma,\rho)=0:\ \gamma\in F_{2}^{m}\backslash\{0\},\ \rho\in F_{2}^{n},\ \forall\ 1\leq wt(\rho)\leq k\}$ (17)
+CI shows variant behavior when an affine transformation is applied to $S$ [88].
+For a good cryptographic S-box the lower is the order of CI the better.
 CI shares an inverse relation with NL, hence if CI holds a lower order the value of NL will be high, and vice versa.
@@
 And these differentials when combined for multiple rounds become differential characteristics of a cipher under study [59].
-The probabilities of occurrence of certain output (ciphertext) differences corresponding fixed input (plaintext) differences are computed in the form of DDT table.
-A low value of differential uniformity, high differential branch number, high value of propagation characteristics, and no undisturbed bits are a requirement for an sbox to resist against differential cryptanalysis and its variant attacks.
+The probabilities of occurrence of certain output (ciphertext) differences corresponding fixed input (plaintext) differences are computed in the form of DDT.
+A low value of differential uniformity, high differential branch number, high value of propagation characteristics, and no undisturbed bits are a requirement for an S-box to resist against differential cryptanalysis and its variant attacks.
 Difference Distribution Table (DDT)
@@
 $\triangle_{\flat}=\flat\oplus\flat^{\prime}$ (19)
-[…]
+Where $\varkappa,\varkappa^{\prime}\in F_{2}^{n}$, $\flat,\flat^{\prime}\in F_{2}^{m}$.
+Ideally, the probability that a particular output difference $(\triangle\flat)$ occurs given a particular input difference $(\triangle\varkappa)$ is $\frac{1}{2^{n-1}}$.
+Same as LAT, the lookup table of DDT is also of the dimension $2^{n}\times 2^{m}$.
+Mathematically:-
 $\mathsf{DDT(\triangle_{\varkappa},\triangle_{\flat})}=\#\{S(\varkappa\oplus\triangle_{\varkappa})\oplus S(\varkappa)=\triangle_{\flat}\}$ (20)
 When an affine transformation is applied to $S$, DDT preserves the maximum entry and weight of coefficients in rows and columns, however their position is subject to a change [117, 118].
-Definition 3.4
+Table 4 illustrates bounds on the possible values a single input difference can map to against a single output difference, except for $\triangle_{\varkappa}=0$.
+Notably, the sum of all entries for a given $\triangle_{\varkappa}$ is upper bounded by $2^{n}$.
+Definition 4.
 The maximum achievable value in the DDT of $S$ is called the differential uniformity (DU) of $S$. It can be mathematically represented as follows [119]:-
 $\mathsf{DU_{S}}=\max_{\triangle_{\varkappa}\neq 0,\triangle_{\flat}}\#\mathsf{DDT(\triangle_{\varkappa},\triangle_{\flat})}$ (21)
-For an sbox to resist differential cryptanalysis a smaller value of DU is required. If an affine transformation is applied to $S$. DU stays invariant [104].
+For an S-box to resist differential cryptanalysis a smaller value of DU is required. If an affine transformation is applied to $S$. DU stays invariant [118].
 Differential Branch Number (DBN)
@@
 $\mathsf{BN_{D}(S)}=\min_{\varkappa,\varkappa^{\prime}\in F_{2}^{n},\varkappa\neq\varkappa^{\prime}}\{wt(\varkappa\oplus\varkappa^{\prime})\oplus wt(S(\varkappa)\oplus S(\varkappa^{\prime}))\}$ (22)
-An sbox having DBN greater than $2$ increases the chances of having more active sboxes in iterative ciphers. DBN does not stay invariant under affine equivalence [98].
+An S-box having DBN greater than $2$ increases the chances of having more active S-boxes in iterative ciphers. DBN does not stay invariant under affine equivalence [112].
 Propagation Criteria (PC)
@@
 $PC(l)$ of $S$ is not variant when an affine transformation is applied [101].
-An sbox satisfies PC when $ACT_{max}=0$ For an sbox to satisfy SAC, complying with PC is a pre-requisite [74].
-Sboxes having low order $PC(l)$ have weak diffusion characteristics [73].
+An S-box satisfies PC when $ACT_{max}=0$.
+For an S-box to satisfy SAC, complying with PC is a pre-requisite [88].
+S-boxes having low order $PC(l)$ have weak diffusion characteristics [87].
 Undisturbed Bits (UDB)
-For a specific nonzero input difference $\triangle_{\varkappa}\in F_{2}^{n}$ of an sbox $S:F_{2}^{n}\rightarrow F_{2}^{m}$, some of the bits in the corresponding output difference $\triangle_{\rho}$ remain unchanged, these specific bits are referred to as undisturbed bits (UDB) [108].
+For a specific nonzero input difference $\triangle_{\varkappa}\in F_{2}^{n}$ of an S-box $S:F_{2}^{n}\rightarrow F_{2}^{m}$, some of the bits in the corresponding output difference $\triangle_{\rho}$ remain unchanged, these specific bits are referred to as undisturbed bits (UDB) [122].
 Mathematically:-
-=$\triangle_{\flat}=\{\triangle_{\rho}=\rho_{j}\in F_{2}^{m}\ |Pr_{S}(\triangle_{\varkappa}\rightarrow\triangle_{\rho})>0\}$ (24)
-Here $j=(0,1,2,....m-1)$, if $\rho_{j}=d$, when $d\in F_{2}$ is fixed and $\triangle_{\rho}\in\triangle_{\flat}$ then $S$ has UDBs.
+$\triangle_{\flat}=\{\triangle_{\rho}=\rho_{j}\in F_{2}^{m}\ |Pr_{S}(\triangle_{\varkappa}\rightarrow\triangle_{\rho})>0\}$ (24)
+Here j=(0,1,2,.m1)j=(0,1,2,....m-1), if $\rho_{j}=d$, when $d\in F_{2}$ is fixed and $\triangle_{\rho}\in\triangle_{\flat}$ then $S$ has UDBs.
 UDB is invariant under affine transformation.
-An sbox should have no UDB in order to avoid the construction of truncated differentials.
+An S-box should have no UDB in order to avoid the construction of truncated differentials.
 Currently, there is no theoretically proven upper bound on UDB.
@@
 However, instead of drawing one differential characteristic for the entire cipher, boomerang attack divides the cipher into two sub ciphers and generates two differentials characteristics also known as boomerang distinguishers.
-A low value of boomerang uniformity makes an sbox resistant towards boomerang attacks.
+A low value of boomerang uniformity makes an S-box resistant towards boomerang attacks.
 Boomerang Connectivity Table (BCT)
@@
 Boomerang Uniformity (BU)
-The highest value in BCT is known as the boomerang uniformity (BU); where $[\triangle_{\varkappa},\nabla_{\flat}]\neq 0$ [55].
+The highest value in BCT is known as the boomerang uniformity (BU); where $\triangle_{\varkappa},\nabla_{\flat}]\neq 0 [$67].
 $\beta\mathsf{U}=\max_{\triangle_{\varkappa},\nabla_{\flat}\in F_{2}^{n},F_{2}^{m}\backslash\{0\}}\beta T(\triangle_{\varkappa},\nabla_{\flat})$ (26)
-BCT and BU of an sbox are invariant under affine transformation [74, 55].
+BCT and BU of an S-box are invariant under affine transformation [88, 67].
 3.5 Differential-Linear Cryptanalysis
@@
 For $\triangle_{\varkappa}\in F_{2}^{n}$ and $\rho\in F_{2}^{m}$, the DLCT entry $(\triangle_{\varkappa},\rho)$ can be computed mathematically as follows:-
-$\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho)=\#\{\varkappa\in F_{2}^{n}|\ \rho\cdot S(\varkappa)=\rho\cdot S(\varkappa+\triangle_{\varkappa})\}-2^{n-1}$ (27)
+$\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho)=\#\{\varkappa\in F_{2}^{n}\mid\ \rho\cdot S(\varkappa)=\rho\cdot S(\varkappa+\triangle_{\varkappa})\}-2^{n-1}$ (27)
 Differential Linear Uniformity (DLU)
@@
 Mathematically:-
-$\Gamma_{S}=\{\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho),\triangle_{\varkappa}\in F_{2}^{n}\backslash\{0\}\mbox{ and }\rho\in F_{2}^{m}\backslash\{0\}\}$ (28)
+[…]
 $\mathsf{DLU}_{S}=\max(\Gamma_{S})$ (29)
-Under affine equivalence, DLCT and DLU are invariant [111, 78]. In order to resist against differential linear attack, an sbox should have low DLU.
+Under affine equivalence, DLCT and DLU are invariant [125, 92]. In order to resist against differential linear attack, an S-box should have low DLU.
 3.6 Algebraic Cryptanalysis
@@
 Unlike the linear component, the number of variables for nonlinear layer increases after every next round.
-Optimal sboxes are better resistant towards algebraic crypltanalysis if they have a high value of algebraic degree and algebraic immunity, and high degree of interpolation polynomial.
+Optimal S-boxes are better resistant towards algebraic crypltanalysis if they have a high value of algebraic degree and algebraic immunity, and high degree of interpolation polynomial.
 Algebraic Degree (AD)
@@
 The AD is an affine invariant property of any mapping $S$ [115, 88].
-Sboxes having a low AD are vulnerable to higher order differential attacks.
+S-boxess having a low AD are vulnerable to higher order differential attacks.
 AD has an inverse relation with CI, the higher is the order of CI, the lower is AD.
@@
 In 1996, Jakobsen and Knudsen introduced interpolation attack on block ciphers [126].
-In this attack the round function or sboxes are represented in the form of algebraic expressions.
+In this attack the round function or S-boxes are represented in the form of algebraic expressions.
 In either case, Lagrange interpolation technique is used to evaluate the algebraic expression.
@@
 Attack complexity depends on number of terms or degree of the resulting polynomial expression.
-For any boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ with input $\varkappa_{i}=\{\varkappa_{1},.....,\varkappa_{n}\}$ and output $\flat_{i}=\{\flat_{1},.....,\flat_{n}\}$ where $i=1,......,n$, we find the Interpolation Polynomial (IP) $P_{i}(\varkappa)$ as follows:
+For any vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ with input $\varkappa_{i}=\{\varkappa_{1},.....,\varkappa_{n}\}$ and output $\flat_{i}=\{\flat_{1},.....,\flat_{n}\}$ where i=1,,ni=1,......,n, we find the Interpolation Polynomial (IP) $P_{i}(\varkappa)$ as follows:
 $P_{i}(\varkappa)=\sum_{i=1}^{n}\flat_{i}\prod_{j=1;j\neq i}^{n}\frac{\varkappa-\varkappa_{j}}{\flat_{i}-\flat_{j}}$ (33)
-IP is invariant under affine equivalence [113], and a low degree of IP makes any sbox susceptible to interpolation attack.
+IP is invariant under affine equivalence [127], and a low degree of IP makes any S-box susceptible to interpolation attack.
 3.7 Side Channel Analysis
@@
 Due to the increasing threat, designers are always trying to come up with efficient ways to resist against side channel attacks [128].
-Starting from the design of components such as sboxes with good cryptographic properties to resist side channels and moving up to the overall cipher level by employing countermeasures such as boolean or polynomial masking.
-Properties such as DPA SNR and transparency order are nowadays actively used to determine an sbox’s resilience against DPA attacks.
+Starting from the design of components such as S-boxes with good cryptographic properties to resist side channels and moving up to the overall cipher level by employing countermeasures such as boolean or polynomial masking.
+Properties such as DPA SNR and transparency order are nowadays actively used to determine an S-box’s resilience against DPA attacks.
 However, striking a perfect balance between the metrics related to traditional cryptanalysis and side channel analysis is an open research area.
 Currently, designers try to achieve cryptographic profiles that are well suited with the requirements of a particular infrastructure or standardization process (as in NIST LWC).
-In ideal terms, for an sbox to be resistant towards attacks such as DPA, it needs to posses a low DPA SNR and transparency order.
+In ideal terms, for an S-box to be resistant towards attacks such as DPA, it needs to posses a low DPA SNR and transparency order.
 DPA SNR
@@
 $\begin{split}\mathsf{TO}=\max_{\beta\in F_{2}^{m}}(\big|m-2wt(\beta)\big|)\times\frac{1}{2^{2n}-2^{n}}\times\\ \left(\sum_{\gamma\in F_{2}^{n}}\bigg|\sum_{\rho\in F_{2}^{m},wt(u)=1}(-1)^{\rho.\beta}\triangle_{S}(\gamma,\rho)\bigg|\right)\end{split}$ (35)
-If an sbox has small value of TO, it is considered highly resistant against DPA attacks [61]. When affine transformation is applied to an sbox, TO does not stay invariant [81].
-4 Results and Analysis
-Table 5 presents the sbox analysis results of Romulus, ASCON/ ISAP, Elephant, Photon-Beetle, and GIFT-COFB. Results of ASCON and ISAP are listed in the same column, since both have same sbox.
-Diffusion Metrics
-If we look at the basic avalanche characteristics of all these sboxes, they satisfy the criteria for balancedness, ensuring equal number of $1$’s and 00’s in their coordinate and component functions.
-This helps reduce the chances of getting any high probability linear approximations, in turn assuring the resistance of these sboxes towards linear cryptanalysis[79].
-The sboxes under analysis are bijective, so they satisfy the criteria of permutation.
-From the bounds discussed in Table 3, we have developed an understanding that ideally the sboxes should have a full cycle length for order of permutation.
-Presumably, sboxes having short iterative periods are vulnerable to attacks so designers take advantage of the affine equivalence criteria and try to improve the order of permutation of sboxes using affine transformation [50, 51, 83, 84].
-Out of the six, Ascon/ ISAP sbox has a FP at $\varkappa=255$.
+If an S-box has small value of TO, it is considered highly resistant against DPA attacks [73]. When affine transformation is applied to an S-box, TO does not stay invariant [95].
+4 Discussion and Results
+We have given the details of our findings on Romulus, ASCON/ ISAP, Elephant, Photon-Beetle, and GIFT-COFB in Tables 6-7 against dedicated cryptographic profiles.
+We briefly explain them in subsequent subsections.
+4.1 Diffusion Metrics
+If we look at the basic avalanche characteristics of all these S-boxes they satisfy the criteria of balancedness.
+This helps reduce the chances of getting any high probability linear approximations, in turn assuring the resistance of these S-boxes towards linear cryptanalysis [93].
+The S-boxes under analysis are bijective, so they satisfy the criteria of permutation.
+From the bounds discussed in Table 4, we have developed an understanding that ideally the S-boxes should have a full cycle length for order of permutation.
+Presumably, S-boxes having short iterative periods are vulnerable to attacks.
+In order to avoid such attacks designers can use affine equivalence to improve the order of permutation of these S-boxes [62, 63, 97, 98].
+Out of the six, Ascon/ ISAP S-box has a FP at $\varkappa=255$.
 Photon-Beetle/ Elephant/ GIFT-COFB, each has one OFPs at $\varkappa=14/8/3$.
-Sboxes having FPs can alter the overall diffusion power that an iterative cipher has to offer and make it vulnerable to invariant subspace attacks, it can be fixed using affine equivalence.
-However, in practice, many notable cipher designs avoid the presence of FPs, due to the reason of achieving good results in other affine variant properties.
-BIC value for all sboxes is $1.0000$ which indicates that the input and output variables have a high correlation amongst them.
-Furthermore, none of sboxes satisfy SAC.
-Both these findings indicate that an attacker can gain statistical information about the sboxes and use it to exploit the entire cryptosystem [85, 115].
-AI values for all the sboxes meet UB, and the values for that of SSI seem a bit better but do not satisfy the LB requirement.
-Hence, results from both the properties suggest that sboxes have weak diffusion characteristics.
-Linear Metrics
-Security towards linear cryptanalysis attacks is ensured when an sbox has a low LAP and high (close to maximum) degree of NL.
-All $6$ sboxes under analysis have obtained the desired bounds on both properties.
-LBN of all $6$ sboxes is closer to the LB, which is opposite to the ideal requirements.
+S-boxes having FPs can alter the overall diffusion power that an iterative cipher has to offer and make it vulnerable to invariant subspace attacks, it may be rectified using affine equivalent S-box.
+The upper bounded BIC value for all the S-boxes indicates that the input and output avalanche vectors are fully correlated.
+Subsequently, none of the S-box satisfies the principles of completeness and avalanche depicted by their SAC failure.
+The AI values for all the S-boxes achieve the upper bound, and the values for that of SSI seem a bit better but do not satisfy the lower bound requirement.
+As discussed earlier, for an S-box to satisfy SAC, one of the prerequisites is compliance with PC.
+In Tables 6-7 we can see that all S-boxes have PC =0=0, which again points towards the weak avalanche profile.
+These findings indicate that an attacker can utilize this information to mount statistical attacks on the underlying cipher [99, 129].
+4.2 Linear Metrics
+Security towards linear cryptanalysis attacks is ensured when an S-box has a low LAP and higher non linearity.
+Our results show that all $6$ S-boxes under analysis have achieved the desired bounds for both properties.
+The LBN of all $6$ S-boxes is closer to the LB, which is opposite to the ideal requirements.
 However, if the design requires it, LBN can be improved using affine transformations.
-The presence of nonzero LS makes it possible for the cipher to leak significant amount of information if specific group(s) of bits in the input are flipped.
-All the sboxes are first order correlation immune, which makes them secure against zero-correlation attacks.
-Differential and Boomerang Metrics
-Almost Perfect Nonlinear (APN) sboxes, in odd dimensions are differentially $2$ uniform [51].
-The only odd dimension case in sboxes under study is of Ascon/ ISAP, and since its has a DU: $8$, we can classify it as not an APN.
-All these sboxes have DU values within the desired bounds.
-DBN of Elephant and Photon-Beetle satisfies the UB, Ascon/ ISAP is closer to the UB and for Romulus and GIFT-COFB DBN value stands on the LB.
-A good DBN ensures more number of active sboxes across multiple rounds in block ciphers.
-As discussed earlier that for an sbox to satisfy SAC one of the pre-requisites is compliance with PC.
-In Table 5 we can see that all sboxes have PC $=0$, which again points towards the weak avalanche profile.
-UDB are detected in all $6$ sboxes.
-In [108] authors have extensively described how the presence of UDB in an sbox impacts other cryptographic properties, such as the occurrence of LS, as we can see in Table 5.
-C.
-Tezcan used these UDB and LS to devise 4 and 5 round impossible, truncated and improbable differentials, and break 5 out of 12 Ascon rounds [116].
-In regards to boomerang cryptanalysis, the BU for all sboxes is attaining the UB, which is opposite to the required bounds in order to be secure against Boomerang attacks.
-Differential-Linear Metrics
+The presence of LS makes it possible for the cipher to leak significant amount of information if specific group(s) of bits in the input are flipped.
+All the S-boxes are first order correlation immune, which makes them secure against zero-correlation attacks.
+4.3 Differential and Boomerang Metrics
+Almost Perfect Nonlinear (APN) S-boxes, in odd dimensions are differentially $2$ uniform [63].
+The only odd dimension case in S-boxes under study is of Ascon/ ISAP, and since its has a DU: $8$.
+The remaining S-boxes have a DU within the desired bounds.
+The DBN of Elephant and Photon-Beetle S-box satisfies the UB, Ascon/ ISAP is closer to the UB, and for Romulus and GIFT-COFB DBN value stands on the LB.
+A good DBN ensures more number of active S-boxes across multiple rounds in block ciphers.
+The UDB are detected in all $6$ S-boxes.
+In [122] authors have extensively described how the presence of UDB in an S-box impacts other cryptographic properties, such as the occurrence of LS, as we can see in Tables 6-7.
+Tezcan et al. used these UDB and LS to devise 4 and 5 round impossible, truncated and improbab differentials, and break 5 out of 12 Ascon rounds [130].
+In regards to boomerang cryptanalysis, the BU for all S-boxes is attaining the UB, which is opposite to the required bounds in order to be secure against Boomerang attacks.
+4.4 Differential-Linear Metrics
 Combined attacks make it easier for an attacker to break a cipher by defining short differential and linear characteristics, as in the case of differential-linear cryptanalysis.
 Hence, attaining the LB on DLU is crucial to gain resistance against these attacks.
-In case of sboxes under analysis, we can observe from Table 5, that the DLU is a little way up from the LB, hence paving a way for differential-linear attack with a reduced complexity.
-Algebraic Metrics
-Previously, we have talked about the inverse relation between CI and AD, and we observed that all $6$ sboxes have a CI $=0$.
-So given the inverse relation, AD of all sboxes should be obtaining the UB, which in our case is true.
-AD determines the complexity of an sbox’s ANF, the higher the degree is the more complex it is to launch an algebraic attack on a cipher.
-Sboxes like those of Ascon/ ISAP having an AD $=2$ are referred to as quadratic and those having AD $=3$ are known as cubic.
-The functions that posses and AD $=1$ are known as affine.
-Side Channel Metrics
-During design and analysis phase, DPA SNR and TO of sbox are the two metrics that can be used to determine a cipher’s resistance towards SCA (DPA) attacks.
-Sboxes that have a high NL and low DU are prone to SCA attacks and have a high resistance towards differential and linear cryptanalysis.
-Similarly, the sboxes having low values of DPA SNR and TO are resistant towards DPA attacks and vulnerable to differential and linear cryptanalysis.
-Hence, finding the ideal balance between the two categories of metrics is crucial to the security of a block cipher.
-Ascon/ ISAP is attaining the almost ideal bounds in terms of NL and DU.
-Hence, the results of TO and DPA SNR, confirm the notion that Ascon/ ISAP sbox is leaning more on the side of security against linear and differential cryptanalysis.
-Rest of the sboxes also seem to be attaining UBs on both of SCA metrics.
-It is important to note here, that sbox is, although major, yet one mathematical component of a cipher.
-Designers can always rely on some other design components to balance out the security aspect if using an optimally secure sbox comes at the cost of performance.
-Table 5: Results of S-box Analysis
-S# Properties Romulus $8\times 8$ Skinny-$128$ Sbox ASCON/ ISAP $5\times 5$ Sbox Elephant $4\times 4$ Sbox GIFT-COFB $4\times 4$ GIFT Sbox Photon-Beetle $4\times 4$ Present Sbox 1 Bijectivity 2 Balancedness 3 Permutation 4 $\mathsf{OP}$ $140$ $26$ $13$ $7$ $9$ 5 $\mathsf{FP}$ $1$ 00 00 00 00 6 $\mathsf{OFP}$ 00 00 $1$ $1$ $1$ 7 $\mathsf{BIC}$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ 8 $\mathsf{SAC}$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ 9 $\barwedge_{S}$ $256$ $32$ $16$ $16$ $256$ 10 $\doublebarwedge_{S}$ $4194304$ $8192$ $1024$ $1024$ $1024$ 11 $\xi$ $0.25000$ $0.25000$ $0.25000$ $0.25000$ $0.25000$ 12 $\mathsf{NL}$ $64$ $8$ $4$ $4$ $4$ 13 $\mathsf{BN_{L}}$ $2$ $3$ $2$ $2$ $2$ 14 $\mathsf{LS}$ $601$ $91$ $9$ $9$ $9$ 15 $\mathsf{CI}$ 00 00 00 00 00 16 $\mathsf{DU}$ $64$ $8$ $4$ $4$ $6$ 17 $\mathsf{BN_{D}}$ $2$ $3$ $3$ $2$ $3$ 18 $\mathsf{PC}$ 00 00 00 00 00 19 $\mathsf{UDB}$ $258$ $35$ $3$ $3$ $6$ 20 $\beta\mathsf{U}$ $16$ $16$ $16$ $16$ $256$ 21 $\mathsf{DLU}$ $\frac{64}{128}=0.5$ $\frac{8}{16}=0.5$ $\frac{4}{8}=0.5$ $\frac{4}{8}=0.5$ $\frac{4}{8}=0.5$ 22 $\mathsf{AD}$ $6$ $2$ $3$ $3$ $3$ 23 $\mathsf{DPASNR}$ $6.312455$ $3.015113$ $2.398501$ $2.128608$ $2.398501$ 24 $\mathsf{TO}$ $7.174510$ $4.258065$ $3.266667$ $3.533333$ $3.466667$
+In case of S-boxes under analysis, we can observe from Tables 6-7, that the DLU is a little way up from the LB, hence paving a way for differential-linear attack with a reduced complexity.
+4.5 Algebraic Metrics
+Previously, we have talked about the inverse relation between CI and AD, and we observed that all $6$ S-boxes have a CI =0=0.
+So given the inverse relation, AD of all S-boxes should be obtaining the UB, which in our case is true.
+AD determines the complexity of an S-box’s ANF, the higher the degree is the more complex it is to launch an algebraic attack on a cipher.
+S-boxes like those of Ascon/ ISAP having an AD =2=2 are referred to as quadratic and those having AD =3=3 are known as cubic.
+The functions that posses and AD =1=1 are known as affine.
+4.6 Side Channel Metrics
+During design and analysis phase, DPA SNR and TO of S-box are the two metrics that can be used to determine a cipher’s resistance towards SCA (DPA) attacks.
+From the conventional cryptanalysis perspective, S-boxes having a high NL and low DU are prone to SCA attacks and have a high resistance towards differential and linear cryptanalysis, and vice versa.
+Hence, finding the ideal balance between the two axes of attack metrics is crucial to the security of a lightweight block cipher.
+Since the S-boxes of LWC finalists vary in size (4, 5 and 8 bits), a direct comparison of the results we obtained of their side-channel metrics is not fair, since the required optimal bounds for each size are different (see Table 4).
+Hence to draw a meaningful comparison, we normalize the DPA SNR and TO using the ideal or worse case bounds as a reference point.
+  • DPA SNR can be normalized using the identity-case bound [72], as follows:
+    $\mathsf{DPA_{SNR}(I)=\sqrt{m}}$ (36)
+    Here, in the case of a balanced S-box $m$ is the number of input/output bits. We now define the normalized DPA SNR as follows:-
+    $\mathsf{DPA_{SNR}(norm)=\frac{{DPA_{SNR}(S)}-{DPA_{SNR}(I)}}{DPA_{SNR}(S)}}$ (37)
+    The results of the suggested normalization are in the range $[0,1]$.
+Here, a value 0 indicates that the S-box under analysis has a DPA SNR equal to that of an identity S-box, which theoretically is the best DPA SNR possible for an S-box to be resistant towards DPA attacks.
+For example, in the case of 8-bit S-box, the identity has a DPA SNR $2.82$ and the DPA SNR of Romulus is $6.312455$.
+The normalized DPA SNR in the 8-bit case is $0.552$, suggesting that Romulus S-box potentially leaks $55.2\%$ more than the identity S-box.
+In the case of 4-bit sbox, the identity has a DPA SNR $2$, while GIFT, PHOTON and Elephant have DPA SNR $2.128608$ and $2.398501$, respectively.
+The normalized DPA SNR in the case of GIFT is $0.064304$, and in the case of PHOTON and Elephant is $0.1992505$, respectively.
+The normalization in the case of 4-bit S-boxes suggests GIFT has the lowest leakage amongst the three 4-bit S-boxes, as it leaks only $6.4304\%$ more than the 4-bit identity S-box.
+  • TO can be normalized using the worse-case (highest leakage) bound [73], as follows:-
+    $\mathsf{{TO}_{highest-leakage}=m}$ (38)
+    Here, in the case of a balanced S-box $m$ is the number of input/output bits. Based on the worse-case bound, we now define the normalized TO as follows:-
+    $\mathsf{TO(norm)=\frac{TO_{highest-leakage}-{TO_{S}}}{TO_{highest-leakage}}}$ (39)
+    The results of this normalization lie in the range [0,1], where a factor 0 depicts that the TO is the same as the worst-case scenario, i.e., the maximum TO value, implying the highest leakage in the case of DPA attacks.
+Any value greater than 0 suggests relative improvement in the TO value and hence reduced leakage.
+For example, in the case of 8-bit S-box, the worse-case TO is $8$ and the TO of Romulus is $7.174510$.
+The normalized TO in the 8-bit case is $0.1031$, suggesting that Romulus S-box potentially leaks $10.31\%$ less than the worse-case S-box.
+5 Exploring Trade-offs in S-box Design
+Based on our analysis in Section 4, several finalist S-boxes exhibit weaknesses in classical cryptographic metrics — such as fixed points, differential uniformity, non-linearity, linear structures, etc.
+So to explore the possibility of improving the cryptographic robustness of S-boxes beyond those used in the NIST LWC finalists, we generated a new $4\times 4$ bit test S-box using the PEIGEN [75] tool for S-box evaluation and generation, as shown in Table 8, using the following constraints:-
+  • Permutation: ✓
+  • Differential Uniformity: 4
+  • Algebraic Degree: 3
+  • Linear Structures: 0
+  • Implementation Cost: 8
+In PEIGEN, cost is a composite metric that captures the implementation efficiency in terms of both hardware and software, focusing on the following traits:-
+  • Gate Equivalent (GE): refers to a 2-input NAND gate unit,
+  • Bitslice Gate: for software implementations (optimized 32-bits),
+  • Circuit depth in terms of latency (depth vs area), and
+  • Multiplicative complexity: minimizing number of non-linear operations (for side channel secure ciphers where masking and other countermeasures are involved).
+Table 8: Proposed 4-bit S-box
+4-bit 0 4 5 1 c 9 6 7 3 a e 8 b 2 f d
+Fixing the implementation cost ensures that the resultant S-box implementation lies within the acceptable practical bounds for lightweight cryptography implementations, while retaining the desired cryptographic profile.
+5.1 Comparison and Analysis
+Table 7, highlights the comparison between the S-box new 4-bit S-box and and the 4-bit S-boxes from GIFT-COFB, Photon-Beetle and Elephant.
+From the classical cryptanalysis perspective all four S-boxes have the same differential uniformity, non-linearity, linear approximation, algebraic degree and correlation immunity.
+A noticeable difference occurs in the properties of the new S-box when compared to the other three, as follows:-
+- The values for BIC and SAC are better.
+- No linear structures and undisturbed bits.
+- Boomerang uniformity has improved.
+Some of the trade-offs that exist in the properties of the new S-box are as follows:-
+- $3$ fixed points are present in the new S-box while the others have null.
+- DPA-SNR value of $2.5786$ is a little higher than for all three finalists, indicating slightly weaker resistance to side-channel (DPA) leakage, but offers optimal resistance against conventional cryptanalysis.
+- TO of $3.4$ is a little higher than that of Elephant ($3.26$), but a little lower than that of PHOTON and GIFT-COFB, indicating a reasonable balance between leakage and classical security.
+The current findings underscore that even when the classical cryptanalysis properties are optimized within practical bounds, striking a perfect balance with the side-channel metrics may not be possible.
+In practice, identifying such balanced S-boxes remains a trial-and-error process that involves testing multiple generation strategies and parameter choices until an acceptable trade-off is found.
+Moreover, it is important to emphasize that the S-box alone does not define the overall security of a lightweight block cipher.
+Resistance to differential and linear cryptanalysis, as well as to side-channel analysis, also depends heavily on other design aspects such as the diffusion layer, round constants, and key schedule, etc.
+A strong S-box is therefore only one building block of a well-rounded cryptographic primitive.
+6 Scope and Limitations
+Our research presents a comprehensive evaluation of the cryptographic properties of S-boxes used in finalist ciphers of the NIST LWC competition.
+For the sake of clarity and accuracy it is important that we acknowledge the scope and limitations of our analysis.
+Firstly, we would like to clarify that our analysis of S-boxes is conducted from a theoretical and metric-based perspective.
+We do not claim to perform any practical side-channel attacks (e.g., DPA, SPA) or cryptanalytic attacks (e.g., differential, linear) on the full/reduced-rounds structure of NIST LWC finalists.
+Strong S-boxes contribute significantly to overall cipher security.
+We have evaluated the S-boxes individually, based on well-established cryptographic metrics listed in Table 3.
+Such compact analysis is a crucial first-step in determining the strength of the confusion layer when designing a new cipher, but it should be noted that full cipher-level cryptanalysis and side-channel analysis requires holistic evaluation – involving linear layers, modes of encryption, masking strategies, implementation environments, and leakage models.
+These are beyond the scope of this study and can be seen as potential future work.
+Therefore, we specifically refrain from extrapolating our S-box component-level findings to make any assertions on the overall security strength of the cipher.
 7 Conclusion
-In this paper, we have thoroughly discussed $25$ plus well known properties of cryptographic sboxes and described their behaviour under affine transformations.
-In addition, we have also provided an insight into the minimum and maximum achievable bounds of these properties and the desirable bounds for an optimal sbox.
-Our aim was to provide a comprehensive walk-through on the taxonomy of cryptographic properties of sboxes for researchers who are new to the field of design and cryptanalysis of symmetric and lightweight ciphers.
-Since the initiation of NIST lightweight cryptography competition, many attacks have been published on different ciphers in the competition.
-Many of these attacks exploit one of the weakness in the substitution layer.
-We hope that our analysis of these sboxes and overview on the bounds, will serve as a pre-requisite profile for cryptanalysts to determine new attack distinguishers and propose countermeasures to make the ciphers secure against these attacks.
+Since the announcement of NIST lightweight cryptography competition, a considerable number of attacks have been published for exploiting the vulnerabilities in the non-linear component of the competitors under consideration.
+This paper presents a detailed cryptographic security analysis of the S-boxes of six finalists, examining the $24$ cryptographic properties that are relevant to established attack methods (see3).
+These six S-boxes exhibit characteristics such as high SAC and BIC values, degree-2 polynomials, and undisturbed bits — that may increase the susceptibility to known differential, linear, and boomerang attacks.
+The existence of degree 2 polynomials is the primary source of linear structures, followed by undisturbed bits.
+Improving the S-box design may, therefore, help improve their overall security margin of these offered by these permutations.
+8 Declarations
+Author Contribution Statement:
+All authors contributed equally to the writing of the manuscript.
+Author Competing Interest statement: Sajid Ali Khan has position on editorial board of Discover Computing, and was not involved in the review or decisions related to this manuscript.
+The other authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Funding Declaration:
+Not applicable
+Ethics and Consent to Participate declarations:
+Not applicable
+Consent for publication Declaration:
+Not applicable
+Ethics, Consent to Participate, and Consent to Publish declarations: Not Applicable
+Availability of Data: The data that support the findings of this study are available from the corresponding author, upon request.
+Preprint: A preprint of this article is available at [131]